### v2

# A Clique Partitioning-Based Algorithm for Graph Compression

Abstract
Reducing the running time of graph algorithms is vital for tackling real-world problems such as shortest paths and matching in large-scale graphs, where path information plays a crucial role. This paper addresses this critical challenge of reducing the running time of graph algorithms by proposing a new graph compression algorithm that partitions the graph into bipartite cliques and uses the partition to obtain a compressed graph having a smaller number of edges while preserving the path information. This compressed graph can then be used as input to other graph algorithms for which path information is essential, leading to a significant reduction of their running time, especially for large, dense graphs. The running time of the proposed algorithm is $O(mn^{\delta})$, where $0\leq\delta\leq 1$, which is better than $O(mn^{\delta}\log^{2}n)$, the running time of the best existing clique partitioning-based graph compression algorithm (the Feder-Motwani (FM) algorithm). Our extensive experimental analysis show that our algorithm achieves a compression ratio of up to $26\%$ greater and executes up to 105.18 times faster than the FM algorithm. In addition, on large graphs with up to 1.05 billion edges, it achieves a compression ratio of up to 3.9, reducing the number of edges up to $74.36\%$. Finally, our tests with a matching algorithm on sufficiently large, dense graphs, demonstrate a reduction in the running time of up to 72.83% when the input is the compressed graph obtained by our algorithm, compared to the case where the input is the original uncompressed graph.
1 Introduction
Graphs are versatile and powerful when used to model and analyze complex real-world problems in various domains, such as social, biological, and communication networks. As the volume of data generated and stored continues to grow, graphs become increasingly large and complex, presenting significant challenges for processing and analyzing data, including prohibitively high processing times and huge storage requirements. To address these challenges, graph compression techniques such as edge compression are employed to decrease the size of a graph while maintaining its essential properties. Edge compression involves shrinking the edge set of a graph, resulting in a compressed graph with fewer edges than the original. During edge compression it is important to preserve the path information that plays a crucial role in graph’s connectivity, i.e., the ability to reach one vertex from another using a sequence of edges. If the path information is preserved during compression, the compressed graphs can be used as input to several other graph algorithms for which the path information is essential such as matching and all-pairs shortest path algorithms, leading to a significant reduction of their running time, especially for large, dense graphs.
However, to benefit from using the compressed graph as input to a graph algorithm, it is critical that the compression algorithm and the obtained compressed graph meet the following requirements: 1) the compression algorithm must have a low running time so that when used as a preprocessing step for other graph algorithms, like matching, leads to a lower overall execution time than in the case of executing the graph algorithms on the original graph; and 2) the compressed graphs are directly usable as input to other graph algorithms or require minimal modifications by these algorithms. These conditions make the design of graph compression algorithms challenging. In this paper, we address this challenge by focusing on bipartite graphs and propose a new Clique Partitioning-based Graph Compression (CPGC) algorithm that compresses bipartite graphs while preserving their path information. A bipartite graph is a graph where the vertices can be divided into two disjoint sets, such that no two vertices within the same set are adjacent. We theoretically and experimentally show that our algorithm improves the running time and the compression rate over the algorithm proposed by Feder and Motwani [9], i.e., the best existing clique partitioning-based graph compression algorithm used to speed-up the execution of other graph algorithms.
1.1 Related Work.
The problem of representing a graph in a compressed form has been extensively explored, with techniques typically falling into two categories: lossless and lossy compression [2]. The lossless compression retains the entirety of graph information, while the lossy compression sacrifices some details during the compression process. As one of the key requirements of our approach is to preserve the path information, it falls into the lossless compression techniques category. The majority of the proposed lossless graph compression techniques focus on specific domains such as web graphs, social networks, and chemistry networks [2]. Those techniques primarily focus on reducing the storage space [4] and/or efficiently addressing specific graph queries, such as graph pattern matching, community query, etc. [1, 3, 5, 8, 13]. Buehrer and Chellapilla [3] achieved high compression ratios and small depths by using bipartite cliques with frequent itemset mining for compressing web graphs. However, while these techniques achieve high compression ratios, the resulting compressed graphs may not be directly usable as input to algorithms like all-pairs shortest paths, often requiring significant modifications or reconstruction overhead. Similarly, Fan et al. [8] achieved significant compression of large graphs by converting stars, cliques, and paths into super-nodes, achieving speedups in query tasks like shortest distance computations. While this technique is compatible with algorithms like matching, obtaining the path information for all nodes may introduce high overhead, increasing the total running time. While compressed graphs tailored for specific query tasks demonstrate efficiency, seamlessly integrating them with certain graph algorithms, such as matching, remains a challenge due to the critical need for comprehensive path information.
To the best of our knowledge the closest work to ours is that of Feder and Motwani [9] who proposed a graph compression technique that preserves the path information of the original graph while retaining the algorithmic properties of the graph. Using the compressed graph obtained by their technique as input to other graph algorithms, such as bipartite matching [11], edge connectivity [12], and vertex connectivity [7], leads to a significant reduction in their running time. The technique of Feder and Motwani [9] is based on partitioning the graph into bipartite cliques (complete bipartite subgraphs), that is, a set of bipartite cliques that partition the edge set of the graph. They provided an algorithm for compressing bipartite graphs based on partitioning the graph into bipartite cliques that runs in $O(mn^{\delta}\log^{2}n)$ time, where $n$ is the number of vertices, $m$ is the number of edges, and $\delta$ is a constant such that $0\leq\delta\leq 1$. In the following section, we present Feder and Motwani’s work as the background for clique partitioning-based graph compression and provide the motivation for our approach.
2 Background and Motivation
2.1 Background.
Feder and Motwani [9] proposed a graph compression algorithm (called FM in the rest of the paper) that involves finding $\delta$-cliques and replacing each of them with a tripartite graph having two of the partitions the same as those of the clique it replaces, and the third partition composed of a single newly added vertex. A $\delta$-clique in a bipartite graph graph $G(U,W,E)$ with $|U|=|W|=n$ and $|E|=m$, is a complete bipartite subgraph with the left partition $U^{\prime}$ of size $\lceil n^{1-\delta}\rceil$ and the right partition $W^{\prime}$ of size ${k}(n,m,\delta)=\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/{m})}}\Big{\rfloor}$, where $\delta$ is a constant such that $0\leq\delta\leq 1$. A tripartite graph $G^{*}=(U,W,Z,E^{*})$ consists of three disjoint sets of vertices $U$, $W$, and $Z$, $(U\cap W=W\cap Z=U\cap Z=\emptyset)$, and a set of edges $E^{*}$, where each edge connects either a vertex in $U$ to a vertex in $W$, or a vertex in $W$ to a vertex in $Z$, or a vertex in $U$ to a vertex in $Z$. Figure 1(a) and 1(c) show an example of a $\delta$-clique with left partition $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{7},u_{8}\}$ and right partition $\{w_{4},w_{5}\}$ in a bipartite graph $G$, and the corresponding tripartite graph that replaces it in the compressed graph $G^{*}$, respectively. The number of edges in the $\delta$-clique is $14=7\times 2$ while the number of edges in the corresponding tripartite graph is $9=7+2$, thus the number of edges in the compressed graph is reduced by 5.
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$
[…]
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$$z_{1}$
Figure 1: (a) Given bipartite graph $G(U,W,E)$; (b) Neighborhood tree of vertex $u_{2}\in U$ that shows the path $\omega$ taken from root to the vertex $w_{j}\in W$ at the leaf and number of edges $d_{u_{2},\omega}$ for each node using the tuple $(\omega,d_{u_{2},\omega})$; and (c) the tripartite graph that replaces the $\delta$-clique with left partition $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{7},u_{8}\}$ and right partition $\{w_{4},w_{5}\}$ in the compressed graph $G^{*}(U,W,Z,E^{*})$.
The FM algorithm iteratively extracts $\delta$-cliques from the input graph $G(U,W,E)$ by selecting $\hat{k}=k(n,m,\delta)$ vertices based on the neighborhood trees associated with each of the vertices $u_{i}\in U$. A neighborhood tree for a vertex $u_{i}\in U$ is a binary tree whose nodes at level $i$ correspond to a partition of $W$ into sets of size $|W|/2^{i}$. Thus, the root node corresponds to the set of all the vertices in $W$, while each leaf node corresponds to a vertex in $W$. Figure 1(b) shows the neighborhood tree for vertex $u_{2}\in U$ which has six neighbour vertices in $W$. The neighborhood trees are described in more detail in Appendix A.1. Once the cliques are extracted, FM compresses the graph by adding a new vertex set $Z$ to the bipartite graph, thus converting it into a tripartite graph $G^{*}$ (Figure 1(c)), where each vertex in $Z$ corresponds to one $\delta$-clique extracted by the algorithm. Each vertex of the left and right partitions of the $\delta$-clique $C_{q}$ is then connected via an edge to a new vertex $z_{q}\in Z$. This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus compressing the graph. The obtained compressed graph $G^{*}$ preserves the path information of the original graph, i.e., the connectivity between the vertices $u_{i}\in U$ and $w_{j}\in W$ of graph $G$. The running time of the FM algorithm is $O(mn^{\delta}\log^{2}n)$. Feder and Motwani [9] also showed that their FM algorithm can be extended to the case of non-bipartite graphs to obtain compression. In Appendix A.1 we provide a more detailed description of the FM algorithm.
2.2 Motivation.
While the FM algorithm provides a solid foundation for lossless graph compression, there are several aspects that can be improved to achieve better compression and efficiency. In this section, we motivate our approach by delineating three critical aspects of a $\delta$-clique-based graph compression technique aimed at improving both compression ratio and execution efficiency.
Vertex selection. For each $\delta$-clique $C_{q}$, the FM algorithm iteratively chooses $\hat{k}$ vertices from $W$ to form the right partition $W_{q}$ of $C_{q}$ using the neighborhood trees associated to each $u_{i}\in U$. The vertices $w_{j}\in W$ are located at the leaf of the neighborhood trees and are selected by computing a path from the root node to the leaf node of the neighborhood trees. The computation of the path depends on the cumulative sum of degrees, $d_{u_{i},\omega}$, $\forall u_{i}\in U$ of each level of the neighborhood trees. Figure 1(b) shows one such path for selecting $w_{5}\in W$ for the right partition $K_{q}$ of $\delta$-clique $C_{q},q=1$. It is important to note that the selection of a vertex is based on the cumulative degrees $d_{u_{i},\omega}$ at every level of the neighborhood trees, i.e., $\sum_{u_{i}\in U}d_{u_{i},\omega}$, which could overlook a vertex $w_{j}\in W$ with high degree, i.e., with high number of neighbors, and thus, lowering the number of common neighbours found for the right partition of the $\delta$-clique. Additionally, to avoid the selection of the same vertex in the following iteration, the neighborhood trees need to be updated, which makes vertex selection a sequential process. Thus, to select vertices that lead to a larger set of common neighbors and to remove the dependency for the vertex selection, our algorithm selects vertices for the right partition of the $\delta$-cliques based on the non-increasing order of their degrees. We now describe the update required for the neighborhood trees to avoid repetitive selection of a vertex in $K_{q}$.
Updating neighborhood trees. With each selected vertex $w_{j}\in W$, the FM algorithm updates each neighborhood tree corresponding to all vertices $u_{i}\in U$ that have an edge with the selected vertex $w_{j}\in W$, i.e., neighborhood trees associated with $U_{K_{q}}=\{u_{i}\in U\ |\ w_{j}\in\Gamma(u_{i})\}$. This update is made by subtracting 1 from $d_{u_{i},\omega}$ at each node on the path from the root to the selected vertex $w_{j}$ (a leaf) regardless of whether the vertices $u_{i}\in U_{K_{q}}$ have an edge with $\forall w_{j}\in K_{q}$. Thus, we can say that with every clique the FM algorithm extracts, it eliminates the selection of at least one vertex $w_{j}\in W$ in future iterations of finding $\delta$-cliques. Thus restricting the extraction of cliques in future iterations even if there exists potential edges that can form non-trivial $\delta$-cliques (i.e., $\hat{k}>1$) improving compression. As we will explain in Section 5, this limitation is particularly relevant in less dense graphs, where there may not be enough common neighbors for the right partition of the $\delta$-clique to form additional cliques using the remaining edges. Therefore, in our approach we choose to update on the degree of each vertex $w_{j}$ in the right partition $K_{q}$ of the $\delta$-clique by subtracting the number of common neighbors found for $K_{q}$. This makes sure that only edges that will be removed by the $\delta$-clique are eliminated for the future iterations to find the $\delta$-clique, thereby giving the vertex $w_{j}\in W$ a chance to be selected for future $\delta$-cliques. We will now discuss the extraction of the $\delta$-cliques.
(a) FM
(b) CPGC
Figure 2: Progression of $\hat{m}$, $\hat{k}$, and number of cliques extracted for a graph with 128 vertices in each bi-partition, density 0.98, and $\delta=1$ by FM and CPGC.
Extracting $\delta$-cliques. A $\delta$-clique is formed by selecting $\hat{k}=k(n,\hat{m},\delta)$ vertices from $W$ in $G$. $\hat{k}$ depends on $\hat{m}$, the number of remaining edges, and on $n$, the number of vertices in $G$. It decreases as the number of iterations of the FM algorithm increases, as shown in Figure 2(a). With each iteration of the algorithm, $|U_{q}|\times|W_{q}|$ edges are removed from $G$ which results in decreasing $\hat{m}$, and thus $\hat{k}$. As more $\delta$-cliques are extracted, both $\hat{m}$ and $\hat{k}$ monotonically decrease. Furthermore, as we observe in Figure 2(a), $\hat{k}$ remains constant for several iterations. This is due to the fact that $\hat{k}$ is computed by using a function involving $\hat{m}$ and $n$, and the algorithm can remove a maximum of $n\times\hat{k}$ edges in each iteration. As a result, $\hat{k}$ is decreased by at most one, leading to a relatively slow decrease. Consequently, $\hat{k}$ remains constant for an increasing number of iterations as the rate of edges being removed decreases with decreasing $\hat{k}$, which motivates the selection of multiple $\delta$-cliques for the same $\hat{k}$ in each iterations of the algorithm. Figure 2(b), shows the progression of $\hat{m}$ and $\hat{k}$ and the number of $\delta$-cliques extracted by our algorithm, CPGC. We observe that for the same $\hat{k}$, unlike FM, CPGC extracts more $\delta$-cliques in one iteration leading to the formation of larger $\delta$-cliques compared to FM. For example, for $\hat{k}=4$, FM extracts one clique in each iteration i.e, from iteration 7 to 11, while CPGC extracts 8 cliques in one iteration. Extracting larger $\delta$-cliques results in removing more edges in each iteration of CPGC and thus accelerates the rate at which $\hat{m}$ decreases. Because of this we also observe that for the same graph CPGC extracted fewer $\delta$-cliques compared to FM. Despite of removing a smaller number of $\delta$-cliques, our experimental results (Section 5) show that CPGC obtains a better compression ratio than FM.
Our contributions. The above limitations of the FM algorithm (the best existing algorithm for clique partitioning-based graph compression) motivated us to design a novel algorithm called Clique Partitioning-based Graph Compression (CPGC). Our contributions are:
  1. 1.
    CPGC enables the removal of multiple $\delta$-cliques in each iteration, thereby obtaining a running time of $O(mn^{\delta})$ which is better than $O(mn^{\delta}\log^{2}n)$, the running time of FM.
  2. 2.
    CPGC enables the selection of vertices for forming cliques more than once. This can increase the number of cliques that are extracted during the compression, particularly in the case of large and high density graphs.
  3. 3.
    CPGC obtains an average compression ratio at least as good as that of the FM algorithm. The compression ratio is defined as $\frac{m}{m^{*}}$, where $m$ is the number of edges in the original graph $G$ and $m^{*}$ is the number of edges in the compressed graph $G^{*}$.
  4. 4.
    A comprehensive experimental analysis showing that on large dense graphs, CPGC achieves a compression ratio of up to 26% greater than that obtained by the FM algorithm and executes up to 105.18 times faster than the FM algorithm.
3 The Proposed Algorithm
3.1 Clique Partitioning-based Graph Compression (CPGC) Algorithm.
In this section, we present the design of our Clique Partitioning-based Graph Compression (CPGC) algorithm. For a given input graph $G$, CPGC builds the compressed graph $G^{*}$ of $G$ by partitioning $G$ into bipartite cliques. CPGC iteratively finds bipartite cliques in the given graph $G(U,W,E)$ and compresses it until finding new bipartite cliques does not contribute to the compression of $G$. The size of the right partition of the cliques is determined by $\hat{k}$, which guarantees a $\delta$-clique, and the selection of vertices $w_{j}\in W$ depends on the degree of vertices $d_{w_{j}}=|N(w_{j})|$. The partition of $G$ into bipartite cliques guarantees that each edge in $G$ is in exactly one clique, i.e., $E(C_{i})\cap E(C_{j})=\emptyset$, $\forall i,j$ and $i\neq j$, where $E(C_{i})$ is the set of edges of $\delta$-clique $C_{i}$. CPGC is given in Algorithm 1. The input of CPGC consists of the adjacency matrix $A$ of $G$, and a constant $\delta$.
Initialization (Line 1). For the given bipartite graph $G(U,W,E)$, CPGC initializes $q$, the index of the bipartite cliques extracted from $G$, the number of vertices, $n$ where $n=|U|=|W|$, the number of edges $\hat{m}$, and the degree of the vertices ${d_{w_{j}}}$ $\forall j=1,\ldots,|W|$, to $[0]_{n}$. CPGC also initializes the set of extracted bipartite cliques $\mathcal{C}$, to the empty set.
Algorithm 1 CPGC: Clique Partitioning-based Graph Compression Algorithm
1:$A$: Adjacency matrix;    $\delta$: a constant such that $0\leq\delta\leq 1$.
2:$q\leftarrow 0;\quad n\leftarrow|W|;\quad\hat{m}\leftarrow|E|;\quad{d_{w}}\leftarrow[0]_{n};\quad\mathcal{C}\leftarrow\emptyset$
3:for ${i=1,\ldots,n}$ do
4:    for ${j=1,\ldots,n}$ do
5:          ${d_{w_{j}}}\leftarrow{d_{w_{j}}}+E_{{u_{i}},w_{j}}$
6:$\hat{k}\leftarrow\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/\hat{m})}}\Big{\rfloor}$
7:while $\hat{k}>1$  do
8:    $\mathscr{C},\hat{A},\hat{d_{w}}\leftarrow\textsf{CSA(}q,\hat{k},A,n,d_{w}\textsf{)}$
9:    $\mathcal{C}\leftarrow\mathcal{C}\cup\mathscr{C}$
10:    $d_{w}\leftarrow\hat{d_{w}}$
11:    $A\leftarrow\hat{A}$
12:    $\hat{m}\leftarrow\sum_{j=1}^{|W|}d_{w_{j}}$
13:    $q\leftarrow q+|\mathscr{C}|$
14:    $\hat{k}\leftarrow\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/\hat{m})}}\Big{\rfloor}$
15:For each clique in $\mathcal{C}=\{C_{1},\ldots,C_{q}\}$ add edges from all the vertices in the left partition $U_{q}$ to an additional vertex $z_{q}$, and from $z_{q}$ to all the vertices in the right partition $K_{q}$. This would compress the graph by replacing $|U_{q}|\times|K_{q}|$ edges with $|U_{q}|+|K_{q}|$ edges.
16:The remaining edges in $G$ are trivial cliques and are added in the compressed graph $G^{*}$.
17:Output: $G^{*}$, the compressed graph of $G$.
Partition size (Lines 2-5). CPGC computes the degree of each vertex $w_{j}\in W$, $d_{w_{j}}=|N(w_{j})|$, where $E_{{u_{i}},w_{j}}$ is 1, if there is an edge between ${u_{i}}$ and $w_{j}$, and 0, otherwise (Lines 2-4). It determines the size of the right partition of the bipartite clique, $\hat{k}$ (Line 5), which guarantees the existence of a $\delta$-clique with $k(n,\hat{m},\delta)$ [9]. It is important to say that $\hat{m}=|E|=\sum_{j=1}^{|W|}d_{w_{j}}$.
Clique extraction (Lines 6-13). CPGC proceeds with extracting $\delta$-cliques until extracting new cliques does not contribute to the graph compression. This happens when $\hat{k}=1$, which results in obtaining trivial bipartite cliques. Thus, the while loop in Lines 6 to 13 is executed until trivial cliques are produced. Clique Stripping Algorithm (CSA), presented in Algorithm 2, finds $\delta$-cliques in the bipartite graph $G$. CSA takes the updated index of the bipartite clique, $q$, the size of the right partition of the bipartite clique, $\hat{k}$, the adjacency matrix, $A$, the size of the left partition of $G$, $n$, and the degree of vertices $d_{w_{j}}=|N(w_{j})|,\ \forall w_{j}\in W$ as input. CSA, which is described in Subsection 3.2, returns the set of bipartite cliques $\mathscr{C}$ that it determined in the current execution, the updated adjacency matrix, $\hat{A}$, and the updated degree of vertices $\hat{d_{w}}$. In Line 8, CPGC adds the new bipartite cliques to the set of all bipartite cliques $\mathcal{C}$, updates the degree of vertices $d_{w}$ (Line 9), and adjacency matrix $A$ (Line 10). In Line 11, CPGC updates the number of edges, $\hat{m}$ in the given graph $G(U,W,E)$ after extracting the $\delta$-cliques, and the clique index $q$ (Line 12). Finally, in Line 13, after the $\delta$-cliques are removed, CPGC updates $\hat{k}$ for the next iterations.
Graph compression (Lines 14-15). When the while loop (Lines 6-13) terminates, CPGC compresses the graph by adding a new vertex set $Z$ to the bipartite graph. Each vertex of the left and right partitions of clique $C_{q}$ are connected to a vertex $z_{q}$ via an edge, thus forming a tripartite graph. This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus compressing the graph. The remaining edges in the given bipartite graph $G(U,W,E)$ connecting the set $U$ and $W$, which are not part of any $\delta$-clique are connected directly in the tripartite graph. Therefore, the compressed graph obtained by the CPGC preserves the path information of the original graph.
In Appendix A.2 we provide an example showing the execution of CPGC on a given bipartite graph.
3.2 Clique Stripping Algorithm (CSA).
The Clique Stripping Algorithm (CSA)., given in Algorithm 2, extracts $\delta$-cliques from the given bipartite graph $G(U,W,E)$. CSA initially selects the vertices for the right partition $K_{c}$ of the clique $C_{c}$ and then the common neighbours for the set $K_{c}$. The input of CSA consists of the updated index of the clique, $q$, the size of the right partition $\hat{k}$ which guarantees a $\delta$-clique in $G$, the adjacency matrix $A$, the size of left/right partition $n$, and the degree of vertices $d_{w_{j}}=|N(w_{j})|$.
Algorithm 2 CSA: Clique Stripping Algorithm
1:$q$: Index of bipartite clique;   $\hat{k}$: Size of the right bipartition in bipartite clique;
2:      $A$: Adjacency matrix;           $n$: Number of vertices in left/right bipartition of $G$;
3:      $d_{w}$: Degree of vertices in bipartition $W$.
4:$\mathcal{K}\leftarrow\emptyset$
5:Sort $d_{w}$ in non-increasing order.
6:Let $d_{w_{\pi(1)}},d_{w_{\pi(2)}},\ldots,d_{w_{\pi(n)}}$ be the order.
7:$j\leftarrow 1$
8:while $d_{w_{\pi(j)}}\geq d_{w_{\pi(\hat{k})}}$ do
9:    $\mathcal{K}\leftarrow\mathcal{K}\cup\{w_{\pi(j)}\}$
10:    $j\leftarrow j+1$
11:$\gamma\leftarrow\big{\lfloor}{\frac{|\mathcal{K}|}{\hat{k}}}\big{\rfloor}$
12:for  $c=q+1,\ldots,q+\gamma$ do
13:    $K_{c}\leftarrow\{w_{\pi_{(j)}}\in\mathcal{K}\ |\ (c-(q+1))\cdot\hat{k}<j\leq(c-q)\cdot\hat{k}\}$
14:    $U_{K_{c}}\leftarrow\{u_{i}\in U|K_{c}\subseteq N(u_{i})\}$
15:    $C_{c}\leftarrow\{(U_{K_{c}},K_{c})\}$
16:    Update $A$, by removing edges $(U_{K_{c}}\times K_{c})\in C_{c}$
17:    Update $d_{w}$ by subtracting $|U_{K_{c}}|$ from each $d_{w_{\pi(j)}}$, where $w_{\pi(j)}\in K_{c}$
18:$\mathscr{C}\leftarrow\{C_{q+1},\ldots,C_{q+\gamma}\}$
19:Output Set of cliques $\mathscr{C},\hat{A},\hat{d_{w}}$
Initialization (Lines 1-3). CSA initializes the set of selected vertices for clique extraction, $\mathcal{K}$ to the empty set. In Line 2, it sorts $d_{w}$ in non-increasing order of the degrees of vertices $w_{j}\in W$, i.e., $d_{w_{\pi(1)}}\geq d_{w_{\pi(2)}}\geq\ldots\geq d_{w_{\pi(n)}}$. Then, in Line 3 it initializes the index $j$ in the list of sorted vertices to 1.
Vertex selection (Lines 4-7). CSA selects the vertices of a clique to be extracted, based on their degrees. Selecting a vertex with higher degree results in a larger set of neighbours that can be part of the selected clique. Therefore, CSA selects the vertices whose degrees are greater than or equal to the $\hat{k}$-th largest degree in the sorted $d_{w}$, i.e., $\mathcal{K}=\{w_{\pi(j)}\ |\ d_{w_{\pi(j)}}\geq d_{w_{\pi(\hat{k})}}\}$. CSA adds the selected vertices to $\mathcal{K}$ and increments the iterator $j$ (Lines 5-6). Thus, the while loop in Lines 4- 6 is executed until all vertices with degrees greater than $d_{w_{\pi(\hat{k})}}$ are chosen. In Line 7, CSA calculates $\gamma$ to partition the set $\mathcal{K}$ such that the size of a partition is less than or equal to $\hat{k}$. Therefore, every time CSA is executed it finds $\gamma$ cliques.
Clique extraction (Lines 8-13). The for loop (Lines 8-9) first partitions $\mathcal{K}$ into subsets $K_{c}$, such that $\bigcup_{c=q+1,\ldots,q+\gamma}K_{c}=\mathcal{K}$ and $K_{q+i}\bigcap K_{q+j}=\emptyset,\ i=1,\ldots,\gamma,j=1,\ldots,\gamma,i\neq j.$ Then it constructs $U_{K_{c}}$ such that each vertex in $U_{K_{c}}$ is part of the set of common neighbors of $K_{c}$ (Line 10). In Line 11, CSA forms clique $C_{c}$ with partitions $U_{K_{c}}$ and $K_{c}$, where $|U_{K_{c}}|\geq\lceil{n^{1-\delta}\rceil}$ and $|K_{c}|=\hat{k}$. After forming the clique $C_{c}$, in Line 12, the algorithm updates the adjacency matrix by removing the edges in the clique $C_{c}$ and in Line 13, updates the degrees of the vertices that are part of the clique by removing the size of left partition from the degree of each vertex $w_{\pi(j)}\in K_{c}$, i.e., $d_{w_{\pi(j)}}=d_{w_{\pi(j)}}-|U_{K_{c}}|,\ \forall\ w_{\pi(j)}\in K_{c},\ c=q+1,\ldots q+\gamma$. Finally, in Line 14, CSA forms $\mathscr{C}$, the set of all the cliques extracted in the current execution, and in Line 15, it returns the set of cliques $\mathscr{C}$, the updated adjacency matrix $\hat{A}$, and the updated degrees of vertices $\hat{d}_{w}$ to CPGC.
Time complexity of CSA. CSA takes $O(n\log n)$ to sort the degrees of vertices $d_{w}$, in Line 2. The while loop in Lines 4-6, takes at most $O(n)$ to select vertices for the $\delta$-cliques. The for loop (Lines 8-13) executes $O(\gamma)$ times, where $1\leq\gamma\leq\frac{n}{\hat{k}}$. In the for loop, Line 9 and 10 takes $O(\hat{k})$ and $O(n\hat{k})$ to find the right and left partition of the $c$-th $\delta$-clique, respectively. CSA takes $O(n\hat{k})$ to remove the edges in the $c$-th $\delta$-clique, in Line 12. It takes $O(\hat{k})$ to update the degrees $d_{w}\in K_{c}$, in Line 13. Therefore, the total running time of CSA is dominated by Lines 10 and 12 and thus CSA takes $O(n\gamma\hat{k})$ time.
Time complexity of CPGC. CPGC takes $O(n^{2})$ to calculate the degrees $d_{w}$ of vertices in $W$ in Lines 2-4. The running time of the while loop (Lines 6-13) is dominated by the running time of the function CSA in Line 7 which is given by Algorithm 2. CSA takes $O(n\gamma\hat{k})$ to remove $\gamma\hat{k}n^{1-\delta}$ edges. Therefore, on average it takes $O(n^{\delta})$ time to remove one edge from the given graph. Thus, the while loop in CPGC takes $O(mn^{\delta})$ time to extract $\delta$-cliques. In the end, to compress the graph, CPGC in Lines 14-15, takes linear time. Therefore, the total running time of CPGC is $O(mn^{\delta})$ which is dominated by the while loop in Lines 6-13.
4 Properties of CPGC
Theorem 4.1
The compressed graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGC preserves the path information of the original graph $G(U,W,E)$.
  • Proof.
    Provided in Appendix A.3.
In the following, we determine a bound on the number of edges in the compressed graph obtained by CPGC. This bound tells us how good the compression achieved by CPGC is. First, we state a theorem from [9] that guarantees the existence of a $\delta$-clique in a bipartite graph. This theorem will be used in the proofs of CPGC’s compression properties.
Theorem 4.2
[…]
Next, we provide a bound on the minimum number of edges $\hat{m}$ of $G$ required to obtain a compression of $G$ into $G^{*}$ by extracting $\delta$-cliques.
Lemma 4.1
Given a graph $G(U,W,E)$, where $n=|U|=|W|$, $m=|E|$, and a constant $\delta$, $10<\delta<$ 1, if $m<2n^{2-\frac{\delta}{2}}$ then extracting $\delta$-cliques and replacing them with tripartite graphs as done in CPGC does not lead to a compression of $G$.
  • Proof.
    Provided in Appendix A.3.
Theorem 4.3
Let $G(U,W,E)$ be any bipartite graph with $|U|=|W|=n$ and $|E|=m>2n^{2-\frac{\delta}{2}}$, where $\delta$ is a constant such that $10<\delta<$ 1. Then, the number of edges in the compressed graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGC is $|E^{*}|=O\left(\frac{m}{k(n,m,\delta)}\right)$.
  • Proof.
    We follow the basic idea of the proof from Theorem 2.4 in [9] and extend it to apply to the CPGC algorithm. We assume that initially the compressed graph $G^{*}$ has 0 edges, i.e., $|E^{*}|=0$, and that edges are added to $E^{*}$ as the algorithm progress. To estimate the number of edges in $E^{*}$, we divide the iterations of CPGC into stages, where each stage $i$ consists of extracting one or more $\delta$-cliques with a fixed $\hat{k}_{i}$. Therefore, the $i^{th}$ stage includes all $\delta$-cliques extracted after the number of edges in $G$ becomes less than $m/2^{i-1}$ for the first time, and before the number of edges in $G$ becomes less than $m/2^{i}$ for the first time. For stage $i$, $\hat{k}$ is always going to be at least $k_{i}=\frac{\delta\log n}{2\log(2^{i}\cdot 2n^{2}/m)}$.
    Assume that a clique $C_{q}$ with right partition $K_{q}\subseteq W$ and left partition $U_{K_{q}}\subseteq U$ is extracted in stage $i$. Extracting $C_{q}$ removes $|K_{q}\times U_{K_{q}}|$ edges from $G$ and adds $|K_{q}|+|U_{K_{q}}|$ edges in $E^{*}$. Therefore, the average number of edges added in $E^{*}$ for each edge removed from $G$ by extracting $C_{q}$ is $\rho=\frac{|K_{q}|+|U_{K_{q}}|}{|K_{q}\times U_{K_{q}}|}=\frac{1}{|K_{q}|}+\frac{1}{|U_{K_{q}}|}$. From the definition of a $\delta$-clique, during stage $i$ we have $|K_{q}|=\hat{k}\geq k_{i}$ and $|U_{K_{q}}|\geq n^{1-\delta}$, and therefore, $\rho\leq\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}$. The total number of edges removed from $G$ in stage $i$ cannot be greater than $2m/2^{i}$, therefore, the number of edges added to $E^{*}$ during stage $i$ is less than or equal to $\rho\frac{2m}{2^{i}}=\left(\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\right)\frac{2m}{2^{i}}$.
    CPGC terminates extracting cliques when the number of remaining edges in $G$ is less than $2n^{2-\frac{\delta}{2}}$. This is to eliminate the extraction of trivial cliques, as shown in Lemma 4.1. Next, we determine an upper bound on the total number of edges $m^{*}$ added by CPGC to $G^{*}$ before it terminates extracting $\delta$-cliques (i.e., when $m<2n^{2-\frac{\delta}{2}}$). Therefore,
    (4.1) $\begin{split}m^{*}\leq\sum_{i=1}^{\big{\lceil}\log\left(\frac{m}{2n^{2-\delta/2}}\right)\big{\rceil}}\bigg{(}\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\bigg{)}\frac{2m}{2^{i}}\\ \leq\sum_{i=1}^{\infty}\bigg{(}\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\bigg{)}\frac{2m}{2^{i}}.\end{split}$
Since $1/(2k_{i})\leq 1/k+i/(\delta\log n)$, we obtain,
(4.2) $\begin{split}m^{*}\leq\frac{4m}{k}\sum_{i=1}^{\infty}\frac{1}{2^{i}}+\frac{4m}{\delta\log n}\sum_{i=1}^{\infty}\frac{i}{2^{i}}+\frac{2m}{n^{1-\delta}}\sum_{i=1}^{\infty}\frac{1}{2^{i}}\\ \leq 2m\left(\frac{2}{k}+\frac{4}{\delta\log n}+\frac{1}{n^{1-\delta}}\right).\end{split}$
Thus, $m^{*}=O\left(\frac{m}{k}\right)$. After CPGC finishes extracting cliques, there are $2n^{2-\delta/2}$ edges remaining in $G$. Those remaining edges are trivial cliques (i.e., single edges) that are added to $G^{*}$. The number of remaining edges $2n^{2-\delta/2}$ is in $O\left(\frac{m}{k}\right)$. Therefore, the number of edges in the compressed graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGC is $|E^{*}|=O\left(\frac{m}{k}\right)$.
Extension to non-bipartite graphs. In Appendix A.4 we describe how CPGC can be extended to compress non-bipartite graphs.
5 Experimental Results
We investigated the performance of CPGC in terms of both running time and compression ratio, and compared it with the FM algorithm proposed by Feder and Motwani [9]. We could not locate any existing implementation of the FM algorithm. Thus, we independently implemented it in C to the best of our ability. However, we encountered a limitation with the FM algorithm regarding large graphs, where the calculations for selecting vertices may exceed the machine representation limits. Consequently, we could only obtain results for bipartite graphs with up to $n=128$ vertices in each bipartition and 16 thousand edges (i.e., small graphs). To ensure a fair comparison between algorithms, we excluded the time spent on extracting trivial cliques ($\hat{k}=1$) when measuring the runtime of the FM algorithm. CPGC addresses this by excluding trivial cliques considering $\hat{k}>1$, as indicated in Line 6 of Algorithm 1. Furthermore, we investigated the performance of our algorithm, CPGC, on large bipartite graphs with up to 32 thousand vertices in each bipartition and approximately 1.05 billion edges. Both CPGC and FM algorithms were implemented in C and experiments were conducted on a Linux system with an AMD EPYC-74F3 processor, 1 CPU core, 3.2GHz, and 128 GB of memory. We used the GCC compiler (version 8.5.0) for compiling and executing the C code.
Figure 3: CPGC: Average compression ratio (top row: (a), (b), and (c)) and running time (bottom row: (d), (e), and (f)) for graphs with 3.36 million to 1.05 billion edges.
For our experiments, we generated bipartite graphs with various densities using the $G(n,p)$ random graph model adapted to bipartite graphs. In this model, an edge is included in the bipartite graph with probability $p$, independently from every other edge. We implemented a Python program to generate instances of bipartite graphs with the number of vertices $|U|=|W|=n$, where $n$ ranges from 32 to 32,768, and densities of 0.8, 0.85, 0.90, 0.95, and 0.98. This generated graphs with up to 1.05 billion edges, considering $p$ as a measure of the density of the generated bipartite graphs. For each bipartite graph with a given density, we created 10 different instances. We then ran each instance for six different values of delta ($\delta$: 0.5, 0.6, 0.7, 0.8, 0.9, 1), presenting the average values and standard deviations of the running time and compression ratio for these 10 instances.
5.1 Results for Large Graphs.
We present the results obtained from testing CPGC on large bipartite graphs, where each bipartition consists of $n=2^{i}$ vertices, with $i$ ranging from 11 to 15, and a number of edges $m$ ranging from approximately 3.36 million to 1.05 billion. Figures 3a-3c show the average compression ratio, while Figures 3d-3f show the average running time for bipartite graphs.
Average compression ratio. The compression ratio for CPGC is calculated as $\frac{m}{m^{*}}$, where $m$ and $m^{*}$ represent the number of edges in the given and compressed graphs, respectively. We observe that, for the same $n$ and $\delta$, the compression ratio increases with density. For instance, with $n=2^{15}$ and $\delta=0.5$, at a density of $0.80$ with $m=214.75$ million, the compression ratio is $1.95$. In contrast, at a density of $0.98$, with $m=1.05$ billion, the compression ratio increases to $2.44$. This is because the probability of finding common neighbours is correlated with the density, that is, increasing the density increases the probability of finding a large set of common neighbours of the right partition of the $\delta$-clique. Finding more common neighbours results in CPGC extracting more edges, and thus increasing the compression ratio. However, when the density is kept constant, the compression ratio is not monotonic and depends on $\delta$ and $n$. For instance in Figure 3a, when $n=2^{13}$ and $m=53.69$ million for density $0.80$, the compression ratio initially increases from 1.85, for $\delta=0.5$, to 2, for $\delta=0.6$, achieving the maximum compression. Yet, further increasing $\delta$ to 1 leads to a reduction in the compression ratio to 1.36. This is because increasing $\delta$ increases the number of vertices in the right partition $K_{q}$ of a $\delta$-clique, as $|K_{q}|=\hat{k}$ and $\hat{k}\propto\delta$. Consequently, this decreases the probability of finding the common neighbors, and therefore for lower density, increasing $\delta$ results in a decrease in the compression ratio. Conversely, as the density increases, the probability of discovering common neighbors also increases. Hence, at a density of 0.98, an increase in $\delta$ correlates with an increase in the compression ratio for all $n$ and $m$ values (Figure 3c). It is important to note that the compression ratio depends on $n$, $m$, and $\delta$. Certain combinations of $n$, $m$, and $\delta$ lead to the highest compression ratio. We leave the investigation of finding the optimal combination of these parameters for future work.
Figure 4: CPGC vs. FM: Compression ratio relative to FM (top row: (a), (b) and (c)) and speedup relative to FM (bottom row: (d), (e) and (f)) for graphs with 819 to 16 thousand edges.
Average running time. We observe that for the same $\delta$ and density, the running time increases as the number of vertices increases which is expected (CPGC running time is $O(mn^{\delta})$). We also observe that for the same $n$ and density, the running time decreases as $\delta$ increases. For instance, with $m=214.75$ million for density = 0.80 and $\delta=0.5$, the running time is 176.27 seconds, whereas for $\delta=1$ it reduces to 117.8 seconds. This is because of the relation between $\delta$ and $\hat{k}$ described earlier, CPGC cannot find a large set of common neighbours which results in a lower compression ratio for lower density. Also, when CPGC does not find an edge with $u_{i}\in U$ for any vertex $w_{j}\in K_{q}$, then it skips to the next vertex in $U$ to check for an edge with $w_{j}\in K_{q}$ thus, decreasing the running time. Therefore, we do not observe an increase in the running time as $\delta$ increases. However, for density = 0.98, the running time decreases as $\delta$ increases, even when the compression ratio increases. For example, with $m=1.05$ billion for density = 0.80 and $\delta=0.5$, the running time is 164.83 seconds, whereas for $\delta=1$ the running time is 150.97 seconds. This is because for higher density and large $\delta$, large $\delta$-cliques are extracted which decreases the total number of iterations performed by CPGC and subsequently reduces the running time.
Figure 5: Dinitz($G$) vs. Dinitz($G^{*}$): Average running time of Dinitz’s algorithm on the original bipartite graph $G$ (labeled Dinitz($G$)) and on the compressed graph $G^{*}$ (labeled Dinitz($G^{*}$)), for a large graph with approximately 32,000 vertices in each bipartition and 214.75 million to 1.05 billion edges corresponding to densities = $0.80,0.90,$ and $0.98$ and different $\delta$.
5.2 Results for Small Graphs.
In this section, we compare the performance of CPGC against FM on small graphs, particularly, on graphs with $n=2^{i}$ vertices in each bipartition, where $i$ ranges from 5 to 7 and the number of edges ranging from 819 to 16,056. However, due to the limitations of FM as previously mentioned, our comparison is constrained to $n=2^{7}$ and $m=16,056$. We compare the performance of the two algorithms using two metrics, relative compression ratio, and relative speedup.
Compression ratio relative to FM. We define the compression ratio relative to FM, as the ratio of compression ratio achieved by CPGC and FM. Figures 4a-4c show the relative compression ratio relative to FM achieved by CPGC in the case of small graphs, for $\delta=0.5,0.6,0.7,0.8,0.9,1$, and densities = 0.80, 0.90, 0.98. In all cases, the relative compression ratio ranges from $1$ to $1.27$, highlighting that the compression ratio achieved by CPGC is at least as that of FM. CPGC achieves a greater compression ratio than FM because it extracts multiple $\delta$-cliques before $\hat{k}$ decreases as shown in Figure 2(b), resulting in removing more edges at higher values of $\hat{k}$ (Section 3.2). We observe that CPGC achieves a compression ratio equal to that of FM for a small graph with $n=32$, $m=819$, density = 0.80, and $\delta=0.5$. This is because a small graph does not require many iterations for extracting $\delta$-cliques, thereby not giving CPGC the opportunity to extract more cliques over the FM algorithm. However, for the same $n$ and $\delta$ when the graph density increases the compression ratio achieved by CPGC also increases. For example, with $n=32$, $m=921$, and $1,003$, the relative compression ratio is 1.03 and 1.25, respectively. As explained before, certain combinations of $n$, $m$, and $\delta$ lead to the highest compression ratio. For instance, when $n=128$, $m=16,056$ for a density of 0.98, we observe that the relative compression ratio initially increases from 1.1 to 1.14 with increasing $\delta$ from 0.5 to 0.7. However, further increasing $\delta$ to 1 results in a decrease in the relative compression ratio to 1.1.
Speedup relative to FM. We define the speedup relative to FM as the ratio of the runing time of FM over CPGC. Figures 4d-f illustrate the speedup achieved by CPGC for bipartite graphs with $n$ = 32, 64, and 128 vertices in each partition, $m$ = 819 to 16 thousand for densities of 0.80, 0.90, and 0.98, and $\delta$ values ranging from 0.5 to 1. The running time of CPGC is consistently lower than that of FM, resulting in speedup by CPGC across all cases. This is primarily due to the selection of vertices in the while loop (Lines 4-6) in CSA, facilitating faster compression by extracting more than one i.e., $\gamma$, $\delta$-cliques in a single iteration, as shown in Figure 2(b). The speedup increases notably with higher $\delta$ and density. For example, with $n=32$ and $m=819$ for density 0.8, CPGC achieves an average speedup of 3.66 for $\delta=0.5$, increasing to 16.28 for $\delta=1$. Similarly, with $n=32$ and $\delta=1$, CPGC achieves an average speedup ranging from 16.28 for $m=819$ with density 0.8 to 20.97 for $m=16$ thousand with density 0.98. It is worth noting that for small graphs, certain combinations of $n$, $m$, and $\delta$ yield higher speedup values. For instance, with $n=128$ and $m=819$ for density 0.8, the speedup increases from 52.81 to 81.97 for $\delta=0.8$, and then slightly decreases to 60.68 for $\delta=1$.
5.3 Speeding-up Matching Algorithms.
Since the compressed representation of the original graph maintains all the path information, it can be used as input to other graph algorithms such as maximum cardinality matching, to speedup their execution. Consider Dinitz’s algorithm [6] for finding the maximum cardinality matching in a bipartite graph, whose running time is in $O(\sqrt{n}m)$. According to Theorem 4.3 the compression ratio of CPGC is $k=\Omega\Big{(}\frac{\delta\log n}{\log(2n^{2}/{m})}\Big{)}$, where $10<\delta<$ 1. This is the same as the compression ratio achieved by FM [9]. Because the number of edges in the compressed graph is reduced by a factor of $k$, using CPGC as preprocessing for Dinitz’s algorithm leads to a total time of $O(\sqrt{n}m/k+mn^{\delta})$. For $\delta<1/2$, the total running time becomes $O(\sqrt{n}m\log_{n}{(\frac{n^{2}}{m}}))$. To the best of our knowledge, this matches the best known asymptotic bound for maximum cardinality bipartite matching algorithms (also achieved by employing FM instead of CPGC and by the bipartite matching algorithm based on push-relabel method [10]). Similar reasoning applies if instead of Dinitz’s algorithm we consider the Hopcroft-Karp algorithm [11].
Figure 5 shows the average running time of Dinitz’s algorithm on the original bipartite graph $G$ with $n=32$ thousand vertices in each bipartition, and $m$ ranging from 214.75 million to 1.05 billion edges for densities = $0.80,0.90,$ and $0.98$ and different $\delta$. It also shows the run time of Dinitz’s algorithm executed on the compressed tripartite graph $G^{*}$, for different values of $\delta$, and densities $=0.80,0.90,$ and $0.98$. We observe that in all cases the Dinitz algorithm applied to the compressed graph $G^{*}$ outperforms its execution on the original graph $G$, achieving a reduction in the running time of up to 72.83%. Even when considering the preprocessing time required to compress the graph $G$, the running time of the Dinitz algorithm for large, dense graphs is still reduced by up to 25.09% with the compressed graph $G^{*}$. This reduction in the running time is directly proportional to the level of compression achieved, where larger compression ratios result in reduced running time. For instance, as discussed in the context of large graphs (Figure 3(a-c)), for a density of $0.98$, CPGC demonstrates an increasing compression ratio as the edge count $m$ reaches 1.05 billion with increasing values of $\delta$. Consequently, the run time of Dinitz’s algorithm on the compressed graph $G^{*}$ also decreases.
6 Conclusion and Future Work
We proposed a new algorithm for graph compression based on partitioning the graph into bipartite cliques. The proposed algorithm runs in time $O(mn^{\delta})$ which improves over the $O(mn^{\delta}\log^{2}n)$ time of the Feder-Motwani (FM) algorithm (the best existing clique partitioning-based graph compression algorithm). Our algorithm achieves an average compression ratio at least as good as that of FM. Experimental results showed that our algorithm achieved a compression ratio of up to 26% greater than that obtained by FM and executes up to 105.18 times faster than FM. We also investigated the reduction in running time of a cardinality matching algorithm (Dinitz’s algorithm) when using the compressed graph instead of the original graph. The results showed that for sufficient large dense graphs, using the compressed graph obtained by CPGC as an input to the matching algorithm leads to a reduction in the running time of up to 72.83% over the matching algorithm using the original graph. In future work, we plan to continue investigating the graph compression problem and propose algorithms that improve over the compression ratio obtained by our proposed algorithm.
A Appendix
A.1 FM algorithm.
In Algorithm 3 we give a high-level description of the FM algorithm. The FM algorithm first constructs neighborhood trees for each vertex $u_{i}$ in $U$, which are labeled binary trees of depth $r$, whose nodes are labeled by a bit string $\omega$, where $0\leq|\omega|\leq r$. The root node of a neighborhood tree contains all the vertices in $W$ and is labeled by the empty string, $\epsilon$. The bit string $\omega$ of a node is obtained by starting at the root and following the path in the tree up to the node and concatenating a 0 or a 1 to $\omega$ each time the path visits the left, or the right child, respectively. The nodes at level $i$ in the tree correspond to a partition of $W$ into sets of size $|W|/2^{i}$. Thus, each leaf node corresponds to a vertex in $W$. Each node in the neighborhood tree of $u_{i}$ stores $d_{i,\omega}$, the number of neighbors of $u_{i}$ in the set $W_{\omega}$ of vertices of $W$ that are associated with the node with label $\omega$, i.e., $d_{i,\omega}=|N(u_{i})\cap W_{\omega}|$, where $N(u_{i})$ is the set of the neighbors of $u_{i}$. In Figure 1(b), we show the neighborhood tree of vertex $u_{2}$ for the bipartite graph with $|U|=|W|=n=8$ given in Figure 1(a). Vertex $w_{3}$ is a leaf and would have $\omega=010$ in all neighborhood trees of the graph.
Algorithm 3 FM Algorithm: High-level description
1:$G(U,W,E)$: Bipartite graph;
2:Initialize: clique index, $q\leftarrow 0$; no. of vertices $n\leftarrow|U|$; and no. of remaining edges $\hat{m}\leftarrow|E|$.
3:Build neighborhood trees for vertices $u_{i}\in U$.
4:while $\hat{m}\geq n^{2-\delta}$ do
5:    $\hat{k}\leftarrow\bigg{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/\hat{m})}}\bigg{\rfloor}$
6:     Starting with $t=1$ and $U_{t}=U$, iterate while $t\leq\hat{k}$ to select $\hat{k}$ vertices based on $c_{0}$ and $c_{1}$ (calculated as in Equation A.1) to form the ordered set $K_{q}=\{y_{1},y_{2},\ldots,y_{t}\}\subseteq W$. At each stage $t$, obtain set $U_{t+1}=\left\{u\in U_{t}\mid y_{1},\ldots,y_{t}\in N(u)\right\}$ which omits vertices from $U$ that are not adjacent to $K_{q}$. And update the corresponding neighborhood trees such that $d_{i,\omega}$ of the nodes along the path from root to the selected vertex is decreased by 1 for each neighborhood tree corresponding to $U_{t}$.
7:     Form the $\delta$-clique $C_{q}=(U_{q},K_{q})$, where, $U_{q}=U_{t}$.
8:    Update $G$ by removing the edges associated with clique $C_{q}$ and update $\hat{m}\leftarrow|E|$.
9:    $q\leftarrow q+1$
10:For each clique in $\mathcal{C}=\{C_{1},\ldots,C_{q}\}$ add edges from all the vertices in the left partition $U_{q}$ to an additional vertex $z_{q}$, and from $z_{q}$ to all the vertices in the right partition $K_{q}$. This would compress the graph by replacing $|U_{q}|\times|K_{q}|$ edges with $|U_{q}|+|K_{q}|$ edges.
11:The remaining edges in $G$ are trivial cliques and are added in the compressed graph $G^{*}$.
12:Output: $G^{*}$, the compressed graph of $G$.
The algorithm computes $\hat{k}$ and then, in Line 5, it performs $\hat{k}$ iterations (indexed by $t$) to select vertices for the right partition of clique $C_{q}$. In each iteration it selects vertices at leaf nodes of the neighborhood trees by following a path based on $d_{i,\omega}$ of child nodes and the number of distinct ordered subsets of the neighbourhood tree at each level. In order to select a path at each level, the algorithm calculates $c_{0}$ and $c_{1}$ as follows:
(A.1) $c_{j}=C\left(U_{t},\mathscr{C}_{t,\omega\cdot j}\right)=\sum_{i:{u_{i}\in U_{t}}}d_{i,\omega\cdot j}\cdot\left(d_{i}-1\right)^{[{k}-t]},\,j=0,1.$
$c_{j}$ counts the number of ordered sets in $\mathscr{C}_{t,\omega}$ whose elements are adjacent to $u_{i}$, where $\mathscr{C}_{t,\omega}$ denotes the collection of all ordered sets $K$ of $W$ of size $k$ in iteration $t$, with $t\leq k$. In each iteration, a vertex in $W$ is added to the right partition of the clique. If $|N(u_{i})|=d_{i}$, then the number of distinct ordered subsets of $N(u_{i})$ of size $k$ is $d_{i}^{[k]}$, where $d_{i}^{[k]}=d_{i}(d_{i}-1)(d_{i}-2)\cdots(d_{i}-k+1)$. Based on $c_{j}$’s, the FM algorithm chooses either the left node (if $c_{0}\geq c_{1}$) or the right node (if $c_{0}<c_{1}$). When the algorithm reaches a leaf node, it selects the corresponding vertex in $W$, updates the set of common neighbors ($U_{t}$), and decreases $d_{i,\omega}$ of each node along the path to the selected vertex by 1, in neighborhood trees corresponding to $U_{t}$. Thus, it guarantees the selection of a new vertex in the following iterations. The algorithm continues this process until it extracts $\hat{k}$ unique vertices to form the set $K_{q}$, the right partition of the $q$-th $\delta$-clique $C_{q}$.
After forming set $K_{q}$, the FM algorithm proceeds to identify the common neighbors of the vertices that are part of set $K_{q}$ in order to construct $U_{q}$, which is the left partition of the $q$-th $\delta$-clique $C_{q}$. The pair $(K_{q},U_{q})$ forms a $\delta$-clique $C_{q}$. Subsequently, the algorithm updates $\hat{m}$, the number of remaining edges in the graph after extracting each $\delta$-clique. This update is done by subtracting the product of the sizes of $K_{q}$ and $U_{q}$ from $\hat{m}$, i.e., $|U_{q}|\times|K_{q}|$. The algorithm then updates $\hat{k}$ considering the updated value of $\hat{m}$ and repeats the procedure to extract more $\delta$-cliques until no more $\delta$-cliques can be found (determined by the condition of the while loop). The edges that are not removed in this process form trivial cliques.
The FM algorithm compresses the graph by adding a new vertex set $Z$ to the bipartite graph, thus converting it into a tripartite graph (Figure 1(c)), where each vertex in $Z$ corresponds to one clique extracted by the algorithm. Each vertex of the left and right partitions of $\delta$-clique $C_{q}$ is then connected via an edge to a new vertex $z_{q}\in Z$, thus forming a tripartite graph. This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus compressing the graph. The compressed graph obtained by the algorithm preserves the path information of the original graph. The running time of the FM algorithm is $O(mn^{\delta}\log^{2}n)$. The FM algorithm can be extended to the case of non-bipartite graphs, where it computes a compression of a graph in time $O(mn^{\delta}\log^{2}n)$, as shown by Feder and Motwani [9].
A.2 Example: CPGC Execution.
We now show how CPGC works on a bipartite graph $G$ with partitions $U$ and $W$, $n=|U|=|W|=8$, and $|E|=54$, shown in Figure 6(a). CPGC (Algorithm 1), first initializes $q=0,n=|W|=|U|=8,\hat{m}=|E|=54,\mathcal{C}=\emptyset$, and $d_{w}=[0]_{n}$. It then calculates the degree of each vertex $d_{w_{j}}$ in Lines 2-4, as shown in Table 1. It calculates $\hat{k}(n,m,\delta)=2$ in Line 5. Thus, the condition for the while loop in Line 6 is met and CPGC calls CSA in Line 7. With the given inputs, CSA sorts $d_{w}$ in non-increasing order. Let $\{w_{4},w_{2},w_{3},w_{5},w_{6},w_{1},w_{7},w_{8}\}$ be the non-increasing order with corresponding degrees given as $d_{w_{\pi(j)}}$ in Table 1.
CPGC step $j$ 1 2 3 4 5 6 7 8
Initialization $d_{w_{j}}$ 6 7 7 8 7 7 6 6
Sorted $d_{w_{j}}$ $d_{w_{\pi(j)}}$ 8 7 7 7 7 6 6 6
After $1^{st}$iteration $d_{w_{j}}$ 6 0 0 1 0 7 6 6
Table 1: Degrees of vertices in partition $W$ at different steps of CPGC.
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$
$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$$z_{1}$$z_{2}$
Figure 6: Clique partitioning: (a) given bipartite graph $G(U,W)$; (b) extracted clique $C_{1}$; (c) extracted clique $C_{2}$; (d) graph with trivial cliques; and (e) compressed tripartite graph $G(U,Z,W)$
In Line 5, CSA forms set $\mathcal{K}$, where $\mathcal{K}=\{w_{4},w_{2},w_{3},w_{5},w_{6}\}$ has all vertices whose degrees are greater than or equal to $d_{w_{\pi(\hat{k})}}=7$. This results in $\gamma=2$ in Line 7, therefore CSA extracts two cliques in Lines 8-13. Since $q$ is initialized to 0 in CPGC, the for loop in Line 8 iterates two times for $c=1$ and $2$. In Line 9, CSA forms $K_{1}=\{w_{4},w_{2}\}$, and in Line 10, it forms the left partition $U_{K_{1}}$ of bipartite clique $\mathcal{C}_{1}$, by selecting the common neighbors of the vertices in $K_{1}$. For example, for set $K_{1}=\{w_{4},w_{2}\}$, $U_{K_{1}}=U-\{u_{7}\}$ as $w_{4}$ and $w_{2}$ have $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{6},u_{8}\}$ as common neighbors. In Line 12, CSA updates the adjacency matrix $A$ by removing the edges of bipartite clique $\mathcal{C}_{1}$ from the original bipartite graph $G$ and finally in Line 13, it updates all $d_{w}$ by subtracting $|U_{K_{1}}|=7$ from both $d_{w_{4}}$ and $d_{w_{2}}$, thus $d_{w_{4}}=1$ and $d_{w_{2}}=0$. Similarly, when $c=2$, $K_{2}=\{w_{3},w_{5}\}$ and $U_{K_{2}}=U-\{u_{6}\}$. CSA updates the adjacency matrix $A$ and $d_{w}$ such that $d_{w_{3}}=d_{w_{5}}=0$. Therefore, in the first execution, CSA, forms two bipartite cliques and removes a total of $(7\times 2)+(7\times 2)=28$ edges from $G$. The updated degrees of the vertices are shown in Table 1. It then forms the set of bipartite cliques extracted in the current execution in Line 14, $\mathscr{C}=\{C_{1},C_{2}\}$. The two cliques are shown in Figures 6(b) and 6(c).
CSA returns $(\mathscr{C},\hat{A},\hat{d}_{w})$ and then CPGC updates the set $\mathscr{C}=\{C_{1},C_{2}\}$, $d_{w}$, the adjacency matrix $A$, and the number of edges $\hat{m}=54-28=26$, in Lines 8-10. In Line 12, CPGC updates $q=2$ and finally, in Line 13 it updates $\hat{k}$ according to the new value of $\hat{m}$ in Line 11 which results in $\hat{k}=1$. Since $\hat{k}=1$, it means CPGC extracts only trivial bipartite cliques (shown in Figure 6(d)) which do not contribute to graph compression. Thus, it does not meet the condition in the while loop (Line 6) and CPGC terminates.
CPGC then compresses the graph by adding two vertices, $z_{1}$, and $z_{2}$, corresponding to the two cliques $C_{1}$ and $C_{2}$, and adds the corresponding edges to form the tripartite graph. The edges in the given graph $G(U,W,E)$, that are not part of any $\delta$-clique are connected directly as shown in Figure 6(e).
A.3 Proofs
Proof of Theorem 4.1.
  • Proof.
    Let’s assume that CPGC extracts only one $\delta$-clique $C_{q}$ from the given graph $G$. In the $\delta$-clique $C_{q}$, the right partition $K_{q}\subseteq W$ is formed by selecting $\hat{k}$ vertices from $W$ and the left partition $U_{q}\subseteq U$ is formed with the common neighbors of $K_{q}$, forming a complete bipartite graph with $E_{q}$ edges. The compressed graph $G^{*}$ is formed with the same left and right partitions $U$ and $W$ of the graph $G$, and a third partition $Z$ which is a set of additional vertices associated with each of the $\delta$-cliques that CPGC extracts. Our main concern is with the edges $(u_{i},w_{j})\in E^{*}$ in the compressed graph $G^{*}$. $E^{*}$ contains the edges that replace the $\delta$-clique $C_{q}$ by adding the additional vertex $z_{q}\in Z$, and the set of edges $\hat{E}=E\setminus E_{q}$ which were not part of the $\delta$-clique $C_{q}$. It can be easily seen that $\hat{E}\subset E$, that is, the edges in $\hat{E}$ are edges in $G$, and the remaining edges $E^{*}\setminus\hat{E}$ in $E^{*}$ are the edges that replace $C_{q}$ in $G^{*}$. Each edge $(u_{i},w_{j})\in E_{q}$, is replaced by two edges, $(u_{i},z_{q})$ and $(z_{q},w_{j})$, where $u_{i}\in U$, $w_{j}\in W$, and $z_{q}\in Z$. Therefore, for each edge $(u_{i},w_{j})\in E_{q}$, there exists a path from $u_{i}$ to $w_{j}$ composed of two edges, $(u_{i},z_{q})$ and $(z_{q},w_{j})$, that passes through the additional vertex $z_{q}\in Z$, thus preserving the path information. This holds true for all the $\delta$-cliques extracted by CPGC.
Proof of Lemma 4.1.
  • Proof.
    In the $\delta$-clique based graph compression performed by CPGC, the compressed graph $G^{*}(U,W,Z,E^{*})$ is obtained from $G$ by replacing $|E_{q}|=|U_{q}\times K_{q}|$ edges in a $\delta$-clique $C_{q}$ with $|E^{*}|=|U_{q}|+|K_{q}|$ edges in $G^{*}$ and adding an additional vertex $z_{q}\in Z$ for each extracted $\delta$-clique $C_{q}$. The size of the right partition of $C_{q}$ is $|K_{q}|={k}=k(n,m,\delta)=\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/m)}}\Big{\rfloor}$. If $m<2n^{2-\frac{\delta}{2}}$ then the size of the right partition $|K_{q}|$ is $k=1$. Therefore, the number of edges in the $\delta$-clique is $|E_{q}|=|U_{q}|\cdot 1=|U_{q}|$. Those edges are replaced by $|E^{*}|=|U_{q}|+\hat{k}=|U_{q}|+1$ edges in $G^{*}$. Thus, replacing the $\delta$-clique in $G$ with $m<2n^{2-\frac{\delta}{2}}$, the number of edges in $G^{*}$ would actually increase by 1, which does not lead to a compression of $G$.
A.4 Extension to non-bipartite graphs.
CPGC can be extended to compress non-bipartite graphs using a similar technique to that described in [9]. Consider a directed graph $H(V,E)$ with $n$ vertices and $m$ edges. Initially, we transform $H$ into a bipartite graph $G\left(L,R,E^{\prime}\right)$, where each vertex $v\in V$ is duplicated into $v_{L}\in L$ and $v_{R}\in R$. For each directed edge $(u,v)$ in $E$, we create a directed edge $\left(u_{L},v_{R}\right)$ in $E^{\prime}$, with the direction from $L$ to $R$. Furthermore, for each vertex $v\in V$ we add a directed edge from $v_{R}$ to $v_{L}$. The proposed compression algorithm is then applied to the bipartite graph $G$, which includes only the edges from $L$ to $R$. Once the compressed graph is computed we add to it the $n$ directed edges from $u_{R}$ to $u_{L}$, corresponding to each vertex $u\in V$. Since the path information of the original graph is preserved by this transformation, it allows different graph algorithms to work on the compressed graph as well. Undirected general graphs can be transformed first into directed graphs by simply replacing each undirected edge $(u,v)$ in the original graph by two directed edges $(u,v)$ and $(v,u)$, and then applying the technique presented above to transform the directed graph into a bipartite graph. The time complexity and the compression ratio of CPGC are the same as in the case of the bipartite graphs. A similar approach for converting a general graph to a bipartite graph and then compressing the obtained bipartite graph was employed in [3, 5].

### v3

@@
-A Clique Partitioning-Based Algorithm for Graph Compression
+Speeding-up Graph Algorithms via Clique Partitioning
 Abstract
 Reducing the running time of graph algorithms is vital for tackling real-world problems such as shortest paths and matching in large-scale graphs, where path information plays a crucial role.
-This paper addresses this critical challenge of reducing the running time of graph algorithms by proposing a new graph compression algorithm that partitions the graph into bipartite cliques and uses the partition to obtain a compressed graph having a smaller number of edges while preserving the path information.
-This compressed graph can then be used as input to other graph algorithms for which path information is essential, leading to a significant reduction of their running time, especially for large, dense graphs.
-The running time of the proposed algorithm is $O(mn^{\delta})$, where $0\leq\delta\leq 1$, which is better than $O(mn^{\delta}\log^{2}n)$, the running time of the best existing clique partitioning-based graph compression algorithm (the Feder-Motwani (FM) algorithm).
-Our extensive experimental analysis show that our algorithm achieves a compression ratio of up to $26\%$ greater and executes up to 105.18 times faster than the FM algorithm.
-In addition, on large graphs with up to 1.05 billion edges, it achieves a compression ratio of up to 3.9, reducing the number of edges up to $74.36\%$.
-Finally, our tests with a matching algorithm on sufficiently large, dense graphs, demonstrate a reduction in the running time of up to 72.83% when the input is the compressed graph obtained by our algorithm, compared to the case where the input is the original uncompressed graph.
+To address this critical challenge, this paper introduces a graph restructuring algorithm that identifies bipartite cliques and replaces them with tripartite graphs.
+This restructuring leads to fewer edges while preserving complete graph path information, enabling the direct application of algorithms like matching and all-pairs shortest paths to achieve significant runtime reductions, especially for large, dense graphs.
+The running time of the proposed algorithm for a graph $G(V,E)$, with $|V|=n$ and $|E|=m$ is $O(mn^{\delta})$, which is better than $O(mn^{\delta}\log^{2}n)$, the running time of the best existing algorithm for speeding-up other graph algorithms (the Feder-Motwani (FM) algorithm), where $0\leq\delta\leq 1$.
+Both the FM algorithm and the proposed algorithm are originally formulated for bipartite graphs, but can also be applied to general directed or undirected graphs.
+Our extensive experimental analysis demonstrates that the proposed algorithm achieves up to 21.26% higher reduction in the number of edges and runs up to 105.18× faster than the FM algorithm.
+On large synthetic graphs with up to 1.05 billion edges, it attains a reduction in the number of edges of up to 74.36%.
+On real-world graphs, it achieves a reduction in the number of edges by up to 46.8%.
+Furthermore, when used as a preprocessing step, our approach yields up to a 2.07× speedup for the matching algorithms on large synthetic graphs, and up to a 1.74× speedup for the All-Pairs Shortest Path algorithms on real-world graphs, when compared to using the given graph as input.
+keywords: graph algorithms, speeding-up graph algorithms, clique partitioning
 1 Introduction
-Graphs are versatile and powerful when used to model and analyze complex real-world problems in various domains, such as social, biological, and communication networks.
-As the volume of data generated and stored continues to grow, graphs become increasingly large and complex, presenting significant challenges for processing and analyzing data, including prohibitively high processing times and huge storage requirements.
-To address these challenges, graph compression techniques such as edge compression are employed to decrease the size of a graph while maintaining its essential properties.
-Edge compression involves shrinking the edge set of a graph, resulting in a compressed graph with fewer edges than the original.
-During edge compression it is important to preserve the path information that plays a crucial role in graph’s connectivity, i.e., the ability to reach one vertex from another using a sequence of edges.
-If the path information is preserved during compression, the compressed graphs can be used as input to several other graph algorithms for which the path information is essential such as matching and all-pairs shortest path algorithms, leading to a significant reduction of their running time, especially for large, dense graphs.
-However, to benefit from using the compressed graph as input to a graph algorithm, it is critical that the compression algorithm and the obtained compressed graph meet the following requirements: 1) the compression algorithm must have a low running time so that when used as a preprocessing step for other graph algorithms, like matching, leads to a lower overall execution time than in the case of executing the graph algorithms on the original graph; and 2) the compressed graphs are directly usable as input to other graph algorithms or require minimal modifications by these algorithms.
-These conditions make the design of graph compression algorithms challenging.
-In this paper, we address this challenge by focusing on bipartite graphs and propose a new Clique Partitioning-based Graph Compression (CPGC) algorithm that compresses bipartite graphs while preserving their path information.
-A bipartite graph is a graph where the vertices can be divided into two disjoint sets, such that no two vertices within the same set are adjacent.
-We theoretically and experimentally show that our algorithm improves the running time and the compression rate over the algorithm proposed by Feder and Motwani [9], i.e., the best existing clique partitioning-based graph compression algorithm used to speed-up the execution of other graph algorithms.
-1.1 Related Work.
-The problem of representing a graph in a compressed form has been extensively explored, with techniques typically falling into two categories: lossless and lossy compression [2].
-The lossless compression retains the entirety of graph information, while the lossy compression sacrifices some details during the compression process.
-As one of the key requirements of our approach is to preserve the path information, it falls into the lossless compression techniques category.
-The majority of the proposed lossless graph compression techniques focus on specific domains such as web graphs, social networks, and chemistry networks [2].
-Those techniques primarily focus on reducing the storage space [4] and/or efficiently addressing specific graph queries, such as graph pattern matching, community query, etc.
-[1, 3, 5, 8, 13].
-Buehrer and Chellapilla [3] achieved high compression ratios and small depths by using bipartite cliques with frequent itemset mining for compressing web graphs.
-However, while these techniques achieve high compression ratios, the resulting compressed graphs may not be directly usable as input to algorithms like all-pairs shortest paths, often requiring significant modifications or reconstruction overhead.
-Similarly, Fan et al.
-[8] achieved significant compression of large graphs by converting stars, cliques, and paths into super-nodes, achieving speedups in query tasks like shortest distance computations.
-While this technique is compatible with algorithms like matching, obtaining the path information for all nodes may introduce high overhead, increasing the total running time.
-While compressed graphs tailored for specific query tasks demonstrate efficiency, seamlessly integrating them with certain graph algorithms, such as matching, remains a challenge due to the critical need for comprehensive path information.
-To the best of our knowledge the closest work to ours is that of Feder and Motwani [9] who proposed a graph compression technique that preserves the path information of the original graph while retaining the algorithmic properties of the graph.
-Using the compressed graph obtained by their technique as input to other graph algorithms, such as bipartite matching [11], edge connectivity [12], and vertex connectivity [7], leads to a significant reduction in their running time.
-The technique of Feder and Motwani [9] is based on partitioning the graph into bipartite cliques (complete bipartite subgraphs), that is, a set of bipartite cliques that partition the edge set of the graph.
-They provided an algorithm for compressing bipartite graphs based on partitioning the graph into bipartite cliques that runs in $O(mn^{\delta}\log^{2}n)$ time, where $n$ is the number of vertices, $m$ is the number of edges, and $\delta$ is a constant such that $0\leq\delta\leq 1$.
-In the following section, we present Feder and Motwani’s work as the background for clique partitioning-based graph compression and provide the motivation for our approach.
+Graph algorithms are essential for analyzing complex real-world networks across diverse domains like social interactions, biological systems, and communication infrastructures.
+However, as these networks grow in size and complexity, the time required to process them using graph algorithms becomes excessively high.
+This paper tackles the critical issue of high computation time in large-size graphs by introducing a graph restructuring approach that speeds up the execution of graph algorithms such as matching and all pairs shortest paths.
+This approach involves transforming the graph into a more computationally suitable structure, preserving its path information, where path information refers to the existence of paths between vertices.
+Specifically, by restructuring the graph to reduce the number of edges, we aim to speeding-up graph algorithms.
+A key requirement for this is the preservation of complete path information during the graph restructuring process.
+Many crucial graph algorithms, such as matching and all-pairs shortest paths (APSP), rely on this global connectivity information to produce correct results.
+However, to effectively speed up the downstream algorithms, the restructuring algorithm itself must be efficient and the resulting structure readily usable.
+Specifically, it must meet two critical conditions: 1) the graph restructuring algorithm must have a low running time to ensure that its use as a preprocessing step leads to an overall reduction in execution time compared to running the algorithms on the original graph; and 2) the restructured graph should be directly usable as input to these algorithms or require minimal modifications.
+These conditions highlight the inherent challenges in designing effective graph restructuring algorithms for speeding-up graph algorithms.
+In this paper, we address these challenges and propose a new Clique Partitioning based Graph Restructuring (CPGR) algorithm that restructures the graph to reduce the number of edges while preserving the path information.
+We theoretically and experimentally show that our algorithm achieves faster running time and greater reduction in the number of edges compared to the algorithm proposed by Feder and Motwani [14], i.e., the best existing algorithm used to speeding-up the execution of graph algorithms.
+1.1 Related Work
+Table 1: Overview of Graph Restructuring Methods
+Technique Goal Mechanism Applications Limitations for Speeding-up Graph        Algorithms
+Sparsification [15, 18, 20] Reduce # of Edges Edge sub-sampling or filtering Graph        Algorithms,       Visualization Accuracy loss, potential disconnection
+Coarsening [6, 7] Reduce # of Vertices Vertex grouping / aggregation Multilevel Methods, GNN Scaling Loss of fine-grained detail, interpolation error
+Hierarchical Decomposition [2, 4, 12, 19] Reduce Problem Complexity Subdivide problem using graph structure Optimization, Modeling Applicability depends on problem structure
+Lossless Compression [1, 2, 3, 5, 10, 13] Reduce Storage Space Efficient encoding of graph data Storage, Querying Compressed Data Query overhead, less effective on irregular graphs
+Clique Partitioning [14], CPGR [this paper] Speeding-up Graph Algorithms Extracting bipartite cliques Graph Algorithms,        Storage
+Graph-structured data are common across domains such as social networks, biology, and the web, often reaching billions of edges.
+To address the computational challenges posed by such scale, graph restructuring techniques have emerged as a key approach for improving the efficiency of graph algorithms through structural transformation.
+These restructuring methods aim to produce graph representations that are enabling fast computation.
+These approaches include sparsification, graph coarsening, hierarchical decomposition, and lossless compression, each offering trade-offs between speed and accuracy as shown in Table 1.
+Graph sparsification reduces the edge count while approximately preserving properties like cuts, distances, or spectral characteristics, enabling faster execution of algorithms such as PageRank, at the cost of precision [15, 18, 20].
+Graph coarsening merges nodes and edges to create a smaller approximation of the original graph, commonly used in multilevel algorithms for partitioning and clustering [6, 7].
+While this improves scalability, it can obscure fine-grained structural details.
+Hierarchical decomposition organizes graphs into nested clusters or layers, facilitating divide-and-conquer approaches in tasks like community detection and multiscale visualization [2, 4, 12, 19].
+The main drawback lies in the overhead of maintaining and traversing hierarchical structures.
+Lossless compression techniques aim to reduce the graph size while preserving exact recoverability, typically using encoding, redundancy elimination, or succinct data structures [1, 2, 3, 5, 10, 13].
+Although accurate, such methods can add significant overhead.
+Despite their advantages, many of these techniques either compromise path-preservation or introduce substantial preprocessing overhead, limiting their applicability to algorithms that require full connectivity information such as matching and APSP.
+A more targeted approach is path-preserving graph restructuring, proposed by Feder and Motwani [14], that partitions the graph into bipartite cliques to speeding-up downstream graph algorithms such as bipartite matching, APSP, and vertex connectivity.
+Their approach guarantees path preservation but suffers from sequential dependencies during clique extraction, which may limit the performance in terms of execution time and reduction in the number of edges.
+Building on this foundation, our work proposes a novel Clique Partitioning-based Graph Restructuring (CPGR) algorithm.
+Unlike Feder and Motwani’s iterative singleton clique extraction, CPGR introduces multiple and larger clique extraction, enabling more efficient restructuring while preserving exact path information.
+This makes CPGR a promising candidate for preprocessing large-scale graphs to accelerate a wide range of graph algorithms.
 2 Background and Motivation
-2.1 Background.
-Feder and Motwani [9] proposed a graph compression algorithm (called FM in the rest of the paper) that involves finding $\delta$-cliques and replacing each of them with a tripartite graph having two of the partitions the same as those of the clique it replaces, and the third partition composed of a single newly added vertex.
-A $\delta$-clique in a bipartite graph graph $G(U,W,E)$ with $|U|=|W|=n$ and $|E|=m$, is a complete bipartite subgraph with the left partition $U^{\prime}$ of size $\lceil n^{1-\delta}\rceil$ and the right partition $W^{\prime}$ of size ${k}(n,m,\delta)=\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/{m})}}\Big{\rfloor}$, where $\delta$ is a constant such that $0\leq\delta\leq 1$.
-A tripartite graph $G^{*}=(U,W,Z,E^{*})$ consists of three disjoint sets of vertices $U$, $W$, and $Z$, $(U\cap W=W\cap Z=U\cap Z=\emptyset)$, and a set of edges $E^{*}$, where each edge connects either a vertex in $U$ to a vertex in $W$, or a vertex in $W$ to a vertex in $Z$, or a vertex in $U$ to a vertex in $Z$.
-Figure 1(a) and 1(c) show an example of a $\delta$-clique with left partition $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{7},u_{8}\}$ and right partition $\{w_{4},w_{5}\}$ in a bipartite graph $G$, and the corresponding tripartite graph that replaces it in the compressed graph $G^{*}$, respectively.
-The number of edges in the $\delta$-clique is $14=7\times 2$ while the number of edges in the corresponding tripartite graph is $9=7+2$, thus the number of edges in the compressed graph is reduced by 5.
-$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$
+2.1 Background
+Figure 1: (a) Given bipartite graph $G(U,W,E)$ and (b) the tripartite graph that replaces the $\delta$-clique with left partition $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{7},u_{8}\}$ and right partition $\{w_{4},w_{5}\}$ in the restructured graph $G^{*}(U,W,Z,E^{*})$.
+Feder and Motwani [14] proposed a graph restructuring algorithm—referred to as FM in this paper—designed for bipartite graphs $G(U,W,E)$, where $|U|=|W|=n$ and $|E|=m$.
+The algorithm identifies complete bipartite subgraphs called $\delta$-cliques with left and right partitions of size $\lceil n^{1-\delta}\rceil$ and $k(n,m,\delta)=\big\lfloor\frac{\delta\log n}{\log(2n^{2}/m)}\big\rfloor$, respectively, for some constant $0\leq\delta\leq 1$.
+Each $\delta$-clique is then replaced by a more compact tripartite graph that introduces a new vertex $z_{q}$, connecting it to all vertices in the clique’s left and right partitions.
+This transformation reduces the number of edges from $|U_{q}|\times|W_{q}|$ to $|U_{q}|+|W_{q}|$ per clique, achieving significant reduction in the number of edges while preserving the original graph’s connectivity between $U$ and $W$.
+For instance, Figures 1a and 1b show an example of a $\delta$-clique with left partition $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{7},u_{8}\}$ and right partition $\{w_{4},w_{5}\}$ in a bipartite graph $G$, and the corresponding tripartite graph that replaces it in the restructured graph $G^{*}$, respectively.
+The number of edges in the $\delta$-clique is $14=7\times 2$ while the number of edges in the corresponding tripartite graph is $9=7+2$, thus the number of edges in the restructured graph is reduced by 5.
+Thus, the restructured graph $G^{*}=(U,W,Z,E^{*})$ consists of three disjoint vertex sets, where $Z$ contains one new vertex per extracted clique, and edges in $E^{*}$ connect vertices across these partitions.
+Although designed for bipartite graphs, the approach can be extended to general graphs as well.
+Further details on the FM algorithm, including the clique extraction process and the use of neighborhood trees, are provided in Appendix A.1.
+2.2 Motivation
+We motivate our approach by highlighting three critical aspects related to extracting and replacing $\delta$-cliques that influence both the running time and the reduction in the number of edges.
+These aspects arise in existing approaches and suggest opportunities for improving how dense substructures are identified and processed.
+Vertex selection.
+A key challenge in identifying $\delta$-cliques for reducing the number of edges lies in selecting vertices for the right partition $W_{q}$ that share a large set of common neighbors in $U$.
+In approaches such as the FM algorithm, vertices for the right partition are selected using a combinatorial function that accounts for vertex degrees across different levels of a neighborhood tree.
+At each level of the neighborhood tree, vertex degrees within the corresponding partitions are aggregated to determine which branch to explore.
+As a result, vertex selection is driven by combined degree contributions across partitions rather than by the individual potential of vertices to form dense bipartite subgraphs with $U$, which may cause vertices with higher degree to be overlooked when forming dense $\delta$-cliques.
+Figure 2: Progression of $\hat{m}$, $\hat{k}$, and number of cliques extracted for a graph with 128 vertices in each bi-partition, density 0.98, and $\delta=1$ by (a) FM and (b) CPGR.
+Extracting $\delta$-cliques.
+The process of forming $\delta$-cliques relies on selecting $\hat{k}=k(n,\hat{m},\delta)$ vertices from the right partition $W$, where $\hat{m}$ is the number of remaining edges in the graph and $n$ is the number of vertices in each partition.
+As the algorithm progresses and more cliques are extracted, the number of edges $\hat{m}$ decreases, which in turn affects the value of $\hat{k}$.
+Due to the stepwise nature of the function $k(n,\hat{m},\delta)$, the value of $\hat{k}$ often remains constant across multiple iterations.
+As illustrated in Figure 2a, this results in a staircase-like progression in which several iterations operate under the same $\hat{k}$ while removing relatively few edges per iteration.
+This behavior suggests that the iterative structure may not fully exploit the potential of each $\hat{k}$ value, particularly during phases in which $\hat{k}$ remains unchanged over multiple iterations.
+Update mechanism.
+After selecting vertices from the right partition, the manner and timing of degree updates play a crucial role in determining which edges remain available for future $\delta$-clique discovery.
+In the existing approach (FM algorithm), degree updates are applied incrementally as candidate vertices are selected, rather than after a $\delta$-clique has been fully determined.
+As a result, degree reductions may be applied to vertices based on partial clique construction, rather than solely on edges that are ultimately removed by an extracted clique.
+Such early updates can reduce the availability of edges that could participate in subsequent $\delta$-cliques, particularly in sparse or moderately dense graphs.
+This observation highlights the importance of update mechanisms that accurately reflect only the edges eliminated by an extracted clique.
+2.3 Our contributions
+This paper introduces a novel algorithm, the Clique Partitioning-based Graph Restructuring (CPGR) algorithm. Our main contributions are summarized as follows:
+  1. 1.
+    We present CPGR, a graph restructuring algorithm that removes multiple $\delta$-cliques within a single iteration, achieving a running time of $O(mn^{\delta})$.
+This improves upon the $O(mn^{\delta}\log^{2}n)$ running time of the FM algorithm.
+  2. 2.
+    CPGR allows vertices to participate in multiple clique extractions during the restructuring process.
+For a fixed $\hat{k}$, the algorithm is designed to extract as many valid $\delta$ cliques as possible in each iteration, which increases the total number of extracted cliques, particularly in large and high density graphs, and leads to a larger number of edges removed per iteration, as illustrated in Figure 2b.
+  3. 3.
+    CPGR achieves an average reduction in the number of edges that is at least as good as the FM algorithm.
+  4. 4.
+    Experimental analysis show that CPGR achieves up to 21% higher reduction in the number of edges and runs up to 105.18× faster than the FM algorithm.
+  5. 5.
+    Experimentally, we show the practical utility of CPGR as a preprocessing step, achieving a speedup of up to 2.07× for matching and 1.74× for APSP on different graph datasets.
+3 The Proposed Algorithm
+3.1 Clique Partitioning-based Graph Restructuring (CPGR) Algorithm
+This section details the design of our Clique Partitioning-based Graph Restructuring (CPGR) algorithm, presented in Algorithm 1.
+For a given input graph $G$, CPGR obtains a restructured graph $G^{*}$ of $G$ by iteratively finding bipartite cliques (i.e., complete bipartite subgraphs) in $G(U,W,E)$ and replacing them with corresponding tripartite graphs until finding new bipartite cliques does not contribute to the reduction in the number of edges of $G$.
+The size of the right partition of the cliques is determined by $\hat{k}=k(n,\hat{m},\delta)$, which guarantees a $\delta$-clique, and the selection of vertices $w_{j}\in W$ depends on the degree of vertices $d_{w_{j}}=|N(w_{j})|$, where $N(w_{j})$ denotes the set of neighbors of vertex $w_{j}$.
+The partition of $G$ into bipartite cliques guarantees that each edge in $G$ is in exactly one clique, i.e., $E(C_{i})\cap E(C_{j})=\emptyset$, $\forall i,j$ and $i\neq j$, where $E(C_{i})$ is the set of edges of $\delta$-clique $C_{i}$.
+The input of CPGR consists of the adjacency matrix $A$ of $G$, and a constant $\delta$.
+Initialization (Lines 1-5).
+For the given bipartite graph $G(U,W,E)$, CPGR initializes the index $q$ of the bipartite cliques extracted from $G$, the number of vertices, $n$ where $n=|U|=|W|$, the number of edges $\hat{m}$, and the degree of the vertices $d_{w_{j}}$} $\forall j=1,\ldots,|W|$, initialized to $0]_{n}$, the zero vector fo size $n$.
+CPGR also initializes the set of extracted bipartite cliques $\mathcal{C}$, to the empty set.
+Algorithm 1 CPGR: Clique Partitioning-based Graph Restructuring Algorithm
+1:AA: Adjacency matrix;
+2:  $\delta$: a constant such that $0\leq\delta\leq 1$.
 […]
-$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$$z_{1}$
-Figure 1: (a) Given bipartite graph $G(U,W,E)$; (b) Neighborhood tree of vertex $u_{2}\in U$ that shows the path $\omega$ taken from root to the vertex $w_{j}\in W$ at the leaf and number of edges $d_{u_{2},\omega}$ for each node using the tuple $(\omega,d_{u_{2},\omega})$; and (c) the tripartite graph that replaces the $\delta$-clique with left partition $\{u_{1},u_{2},u_{3},u_{4},u_{5},u_{7},u_{8}\}$ and right partition $\{w_{4},w_{5}\}$ in the compressed graph $G^{*}(U,W,Z,E^{*})$.
-The FM algorithm iteratively extracts $\delta$-cliques from the input graph $G(U,W,E)$ by selecting $\hat{k}=k(n,m,\delta)$ vertices based on the neighborhood trees associated with each of the vertices $u_{i}\in U$.
-A neighborhood tree for a vertex $u_{i}\in U$ is a binary tree whose nodes at level $i$ correspond to a partition of $W$ into sets of size $|W|/2^{i}$.
-Thus, the root node corresponds to the set of all the vertices in $W$, while each leaf node corresponds to a vertex in $W$.
-Figure 1(b) shows the neighborhood tree for vertex $u_{2}\in U$ which has six neighbour vertices in $W$.
-The neighborhood trees are described in more detail in Appendix A.1.
-Once the cliques are extracted, FM compresses the graph by adding a new vertex set $Z$ to the bipartite graph, thus converting it into a tripartite graph $G^{*}$ (Figure 1(c)), where each vertex in $Z$ corresponds to one $\delta$-clique extracted by the algorithm.
-Each vertex of the left and right partitions of the $\delta$-clique $C_{q}$ is then connected via an edge to a new vertex $z_{q}\in Z$.
-This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus compressing the graph.
-The obtained compressed graph $G^{*}$ preserves the path information of the original graph, i.e., the connectivity between the vertices $u_{i}\in U$ and $w_{j}\in W$ of graph $G$.
-The running time of the FM algorithm is $O(mn^{\delta}\log^{2}n)$.
-Feder and Motwani [9] also showed that their FM algorithm can be extended to the case of non-bipartite graphs to obtain compression.
-In Appendix A.1 we provide a more detailed description of the FM algorithm.
-2.2 Motivation.
-While the FM algorithm provides a solid foundation for lossless graph compression, there are several aspects that can be improved to achieve better compression and efficiency.
-In this section, we motivate our approach by delineating three critical aspects of a $\delta$-clique-based graph compression technique aimed at improving both compression ratio and execution efficiency.
-Vertex selection.
-For each $\delta$-clique $C_{q}$, the FM algorithm iteratively chooses $\hat{k}$ vertices from $W$ to form the right partition $W_{q}$ of $C_{q}$ using the neighborhood trees associated to each $u_{i}\in U$.
-The vertices $w_{j}\in W$ are located at the leaf of the neighborhood trees and are selected by computing a path from the root node to the leaf node of the neighborhood trees.
-The computation of the path depends on the cumulative sum of degrees, $d_{u_{i},\omega}$, $\forall u_{i}\in U$ of each level of the neighborhood trees.
-Figure 1(b) shows one such path for selecting $w_{5}\in W$ for the right partition $K_{q}$ of $\delta$-clique $C_{q},q=1$.
-It is important to note that the selection of a vertex is based on the cumulative degrees $d_{u_{i},\omega}$ at every level of the neighborhood trees, i.e., $\sum_{u_{i}\in U}d_{u_{i},\omega}$, which could overlook a vertex $w_{j}\in W$ with high degree, i.e., with high number of neighbors, and thus, lowering the number of common neighbours found for the right partition of the $\delta$-clique.
-Additionally, to avoid the selection of the same vertex in the following iteration, the neighborhood trees need to be updated, which makes vertex selection a sequential process.
-Thus, to select vertices that lead to a larger set of common neighbors and to remove the dependency for the vertex selection, our algorithm selects vertices for the right partition of the $\delta$-cliques based on the non-increasing order of their degrees.
-We now describe the update required for the neighborhood trees to avoid repetitive selection of a vertex in $K_{q}$.
-Updating neighborhood trees.
-With each selected vertex $w_{j}\in W$, the FM algorithm updates each neighborhood tree corresponding to all vertices $u_{i}\in U$ that have an edge with the selected vertex $w_{j}\in W$, i.e., neighborhood trees associated with $U_{K_{q}}=\{u_{i}\in U\ |\ w_{j}\in\Gamma(u_{i})\}$.
-This update is made by subtracting 1 from $d_{u_{i},\omega}$ at each node on the path from the root to the selected vertex $w_{j}$ (a leaf) regardless of whether the vertices $u_{i}\in U_{K_{q}}$ have an edge with $\forall w_{j}\in K_{q}$.
-Thus, we can say that with every clique the FM algorithm extracts, it eliminates the selection of at least one vertex $w_{j}\in W$ in future iterations of finding $\delta$-cliques.
-Thus restricting the extraction of cliques in future iterations even if there exists potential edges that can form non-trivial $\delta$-cliques (i.e., $\hat{k}>1$) improving compression.
-As we will explain in Section 5, this limitation is particularly relevant in less dense graphs, where there may not be enough common neighbors for the right partition of the $\delta$-clique to form additional cliques using the remaining edges.
-Therefore, in our approach we choose to update on the degree of each vertex $w_{j}$ in the right partition $K_{q}$ of the $\delta$-clique by subtracting the number of common neighbors found for $K_{q}$.
-This makes sure that only edges that will be removed by the $\delta$-clique are eliminated for the future iterations to find the $\delta$-clique, thereby giving the vertex $w_{j}\in W$ a chance to be selected for future $\delta$-cliques.
-We will now discuss the extraction of the $\delta$-cliques.
-(a) FM
-(b) CPGC
-Figure 2: Progression of $\hat{m}$, $\hat{k}$, and number of cliques extracted for a graph with 128 vertices in each bi-partition, density 0.98, and $\delta=1$ by FM and CPGC.
-Extracting $\delta$-cliques.
-A $\delta$-clique is formed by selecting $\hat{k}=k(n,\hat{m},\delta)$ vertices from $W$ in $G$.
-$\hat{k}$ depends on $\hat{m}$, the number of remaining edges, and on $n$, the number of vertices in $G$.
-It decreases as the number of iterations of the FM algorithm increases, as shown in Figure 2(a).
-With each iteration of the algorithm, $|U_{q}|\times|W_{q}|$ edges are removed from $G$ which results in decreasing $\hat{m}$, and thus $\hat{k}$.
-As more $\delta$-cliques are extracted, both $\hat{m}$ and $\hat{k}$ monotonically decrease.
-Furthermore, as we observe in Figure 2(a), $\hat{k}$ remains constant for several iterations.
-This is due to the fact that $\hat{k}$ is computed by using a function involving $\hat{m}$ and $n$, and the algorithm can remove a maximum of $n\times\hat{k}$ edges in each iteration.
-As a result, $\hat{k}$ is decreased by at most one, leading to a relatively slow decrease.
-Consequently, $\hat{k}$ remains constant for an increasing number of iterations as the rate of edges being removed decreases with decreasing $\hat{k}$, which motivates the selection of multiple $\delta$-cliques for the same $\hat{k}$ in each iterations of the algorithm.
-Figure 2(b), shows the progression of $\hat{m}$ and $\hat{k}$ and the number of $\delta$-cliques extracted by our algorithm, CPGC.
-We observe that for the same $\hat{k}$, unlike FM, CPGC extracts more $\delta$-cliques in one iteration leading to the formation of larger $\delta$-cliques compared to FM.
-For example, for $\hat{k}=4$, FM extracts one clique in each iteration i.e, from iteration 7 to 11, while CPGC extracts 8 cliques in one iteration.
-Extracting larger $\delta$-cliques results in removing more edges in each iteration of CPGC and thus accelerates the rate at which $\hat{m}$ decreases.
-Because of this we also observe that for the same graph CPGC extracted fewer $\delta$-cliques compared to FM.
-Despite of removing a smaller number of $\delta$-cliques, our experimental results (Section 5) show that CPGC obtains a better compression ratio than FM.
-Our contributions.
-The above limitations of the FM algorithm (the best existing algorithm for clique partitioning-based graph compression) motivated us to design a novel algorithm called Clique Partitioning-based Graph Compression (CPGC).
-Our contributions are:
-  1. 1.
-    CPGC enables the removal of multiple $\delta$-cliques in each iteration, thereby obtaining a running time of $O(mn^{\delta})$ which is better than $O(mn^{\delta}\log^{2}n)$, the running time of FM.
-  2. 2.
-    CPGC enables the selection of vertices for forming cliques more than once.
-This can increase the number of cliques that are extracted during the compression, particularly in the case of large and high density graphs.
-  3. 3.
-    CPGC obtains an average compression ratio at least as good as that of the FM algorithm.
-The compression ratio is defined as $\frac{m}{m^{*}}$, where $m$ is the number of edges in the original graph $G$ and $m^{*}$ is the number of edges in the compressed graph $G^{*}$.
-  4. 4.
-    A comprehensive experimental analysis showing that on large dense graphs, CPGC achieves a compression ratio of up to 26% greater than that obtained by the FM algorithm and executes up to 105.18 times faster than the FM algorithm.
-3 The Proposed Algorithm
-3.1 Clique Partitioning-based Graph Compression (CPGC) Algorithm.
-In this section, we present the design of our Clique Partitioning-based Graph Compression (CPGC) algorithm.
-For a given input graph $G$, CPGC builds the compressed graph $G^{*}$ of $G$ by partitioning $G$ into bipartite cliques.
-CPGC iteratively finds bipartite cliques in the given graph $G(U,W,E)$ and compresses it until finding new bipartite cliques does not contribute to the compression of $G$.
-The size of the right partition of the cliques is determined by $\hat{k}$, which guarantees a $\delta$-clique, and the selection of vertices $w_{j}\in W$ depends on the degree of vertices $d_{w_{j}}=|N(w_{j})|$.
-The partition of $G$ into bipartite cliques guarantees that each edge in $G$ is in exactly one clique, i.e., $E(C_{i})\cap E(C_{j})=\emptyset$, $\forall i,j$ and $i\neq j$, where $E(C_{i})$ is the set of edges of $\delta$-clique $C_{i}$.
-CPGC is given in Algorithm 1.
-The input of CPGC consists of the adjacency matrix $A$ of $G$, and a constant $\delta$.
-Initialization (Line 1).
-For the given bipartite graph $G(U,W,E)$, CPGC initializes $q$, the index of the bipartite cliques extracted from $G$, the number of vertices, $n$ where $n=|U|=|W|$, the number of edges $\hat{m}$, and the degree of the vertices ${d_{w_{j}}}$ $\forall j=1,\ldots,|W|$, to $[0]_{n}$.
-CPGC also initializes the set of extracted bipartite cliques $\mathcal{C}$, to the empty set.
-Algorithm 1 CPGC: Clique Partitioning-based Graph Compression Algorithm
-1:$A$: Adjacency matrix;    $\delta$: a constant such that $0\leq\delta\leq 1$.
-2:$q\leftarrow 0;\quad n\leftarrow|W|;\quad\hat{m}\leftarrow|E|;\quad{d_{w}}\leftarrow[0]_{n};\quad\mathcal{C}\leftarrow\emptyset$
-3:for ${i=1,\ldots,n}$ do
-4:    for ${j=1,\ldots,n}$ do
-5:          ${d_{w_{j}}}\leftarrow{d_{w_{j}}}+E_{{u_{i}},w_{j}}$
-6:$\hat{k}\leftarrow\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/\hat{m})}}\Big{\rfloor}$
-7:while $\hat{k}>1$  do
-8:    $\mathscr{C},\hat{A},\hat{d_{w}}\leftarrow\textsf{CSA(}q,\hat{k},A,n,d_{w}\textsf{)}$
-9:    $\mathcal{C}\leftarrow\mathcal{C}\cup\mathscr{C}$
-10:    $d_{w}\leftarrow\hat{d_{w}}$
-11:    $A\leftarrow\hat{A}$
-12:    $\hat{m}\leftarrow\sum_{j=1}^{|W|}d_{w_{j}}$
-13:    $q\leftarrow q+|\mathscr{C}|$
-14:    $\hat{k}\leftarrow\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/\hat{m})}}\Big{\rfloor}$
-15:For each clique in $\mathcal{C}=\{C_{1},\ldots,C_{q}\}$ add edges from all the vertices in the left partition $U_{q}$ to an additional vertex $z_{q}$, and from $z_{q}$ to all the vertices in the right partition $K_{q}$.
-This would compress the graph by replacing $|U_{q}|\times|K_{q}|$ edges with $|U_{q}|+|K_{q}|$ edges.
-16:The remaining edges in $G$ are trivial cliques and are added in the compressed graph $G^{*}$.
-17:Output: $G^{*}$, the compressed graph of $G$.
-Partition size (Lines 2-5).
-CPGC computes the degree of each vertex $w_{j}\in W$, $d_{w_{j}}=|N(w_{j})|$, where $E_{{u_{i}},w_{j}}$ is 1, if there is an edge between ${u_{i}}$ and $w_{j}$, and 0, otherwise (Lines 2-4).
-It determines the size of the right partition of the bipartite clique, $\hat{k}$ (Line 5), which guarantees the existence of a $\delta$-clique with $k(n,\hat{m},\delta)$ [9].
+[…]
+[…]
+[…]
+[…]
+8:for $i=1,\ldots,n$ do
+9:  for $j=1,\ldots,n$ do
+10:    ${d_{w_{j}}}\leftarrow{d_{w_{j}}}+E_{{u_{i}},w_{j}}$
+[…]
+12:while $\hat{k}>1$ do
+13:  ($\mathscr{C},\hat{A},\hat{d_{w}})\leftarrow\textsf{CSA(}q,\hat{k},A,n,d_{w}\textsf{)}$
+14:  $\mathcal{C}\leftarrow\mathcal{C}\cup\mathscr{C}$
+15:  $d_{w}\leftarrow\hat{d_{w}}$
+16:  $A\leftarrow\hat{A}$
+17:  $\hat{m}\leftarrow\sum_{j=1}^{|W|}d_{w_{j}}$
+18:  $q\leftarrow q+|\mathscr{C}|$
+19:  $\hat{k}\leftarrow\Big\lfloor{\frac{\delta\log n}{\log(2n^{2}/\hat{m})}}\Big\rfloor$
+[…]
+21:for $i\in\mathcal{C}=\{\mathscr{C}_{1}(U_{1},W_{1}),\ldots,\mathscr{C}_{q}(U_{q},W_{q})\}$ do
+22:  $Z\leftarrow Z\cup\{z_{i}\}$
+23:  $E^{*}\leftarrow E^{*}\cup\{(u,z_{i}),(z_{i},w)\}\quad\forall u\in U_{i},\quad\forall w\in W_{i}$
+24:  $\mathcal{E}_{\mathcal{C}}\leftarrow\mathcal{E}_{\mathcal{C}}\cup E_{i}$
+[…]
+26:Output: $G^{*}(U,Z,W,E^{*})$, the restructured graph of $G$.
+Partition size (Lines 6-9).
+CPGR computes the degree of each vertex $w_{j}\in W$, $d_{w_{j}}=|N(w_{j})|$, where $E_{{u_{i}},w_{j}}$ is 1, if there is an edge between $u_{i}$} and $w_{j}$, and 0, otherwise (Lines 6-8).
+It determines the size of the right partition of the bipartite clique, $\hat{k}$ (Line 9), which guarantees the existence of a $\delta$-clique with $k(n,\hat{m},\delta)$ [14].
 It is important to say that $\hat{m}=|E|=\sum_{j=1}^{|W|}d_{w_{j}}$.
-Clique extraction (Lines 6-13).
-CPGC proceeds with extracting $\delta$-cliques until extracting new cliques does not contribute to the graph compression.
+Clique extraction (Lines 10-17).
+CPGR proceeds with extracting $\delta$-cliques until extracting new cliques does not contribute to the reduction in the number of edges.
 This happens when $\hat{k}=1$, which results in obtaining trivial bipartite cliques.
-Thus, the while loop in Lines 6 to 13 is executed until trivial cliques are produced.
+Thus, the while loop in Lines 10 to 17 is executed until trivial cliques are produced.
 Clique Stripping Algorithm (CSA), presented in Algorithm 2, finds $\delta$-cliques in the bipartite graph $G$.
-CSA takes the updated index of the bipartite clique, $q$, the size of the right partition of the bipartite clique, $\hat{k}$, the adjacency matrix, $A$, the size of the left partition of $G$, $n$, and the degree of vertices $d_{w_{j}}=|N(w_{j})|,\ \forall w_{j}\in W$ as input.
-CSA, which is described in Subsection 3.2, returns the set of bipartite cliques $\mathscr{C}$ that it determined in the current execution, the updated adjacency matrix, $\hat{A}$, and the updated degree of vertices $\hat{d_{w}}$.
-In Line 8, CPGC adds the new bipartite cliques to the set of all bipartite cliques $\mathcal{C}$, updates the degree of vertices $d_{w}$ (Line 9), and adjacency matrix $A$ (Line 10).
-In Line 11, CPGC updates the number of edges, $\hat{m}$ in the given graph $G(U,W,E)$ after extracting the $\delta$-cliques, and the clique index $q$ (Line 12).
-Finally, in Line 13, after the $\delta$-cliques are removed, CPGC updates $\hat{k}$ for the next iterations.
-Graph compression (Lines 14-15).
-When the while loop (Lines 6-13) terminates, CPGC compresses the graph by adding a new vertex set $Z$ to the bipartite graph.
-Each vertex of the left and right partitions of clique $C_{q}$ are connected to a vertex $z_{q}$ via an edge, thus forming a tripartite graph.
-This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus compressing the graph.
-The remaining edges in the given bipartite graph $G(U,W,E)$ connecting the set $U$ and $W$, which are not part of any $\delta$-clique are connected directly in the tripartite graph.
-Therefore, the compressed graph obtained by the CPGC preserves the path information of the original graph.
-In Appendix A.2 we provide an example showing the execution of CPGC on a given bipartite graph.
-3.2 Clique Stripping Algorithm (CSA).
-The Clique Stripping Algorithm (CSA)., given in Algorithm 2, extracts $\delta$-cliques from the given bipartite graph $G(U,W,E)$.
-CSA initially selects the vertices for the right partition $K_{c}$ of the clique $C_{c}$ and then the common neighbours for the set $K_{c}$.
-The input of CSA consists of the updated index of the clique, $q$, the size of the right partition $\hat{k}$ which guarantees a $\delta$-clique in $G$, the adjacency matrix $A$, the size of left/right partition $n$, and the degree of vertices $d_{w_{j}}=|N(w_{j})|$.
+CSA takes the updated index $q$ of the bipartite clique, the size $\hat{k}$ of the right partition of the bipartite clique, the adjacency matrix $A$, the size $n$ of the left partition of $G$, and the degree $d_{w_{j}}=|N(w_{j})|,\ \forall w_{j}\in W$ as input.
+CSA, which is described in Subsection 3.2, returns the set $\mathscr{C}$ of bipartite cliques that it determined in the current execution, the updated adjacency matrix $\hat{A}$, and the updated degree of vertices $\hat{d_{w}}$.
+n Line 12, CPGR adds the new bipartite cliques to the set of all bipartite cliques $\mathcal{C}$, updates the degree of vertices $d_{w}$ (Line 13), and adjacency matrix $A$ (Line 14).
+In Line 15, CPGR updates the number of edges, $\hat{m}$ in the given graph $G(U,W,E)$ after extracting the $\delta$-cliques, and the clique index $q$ (Line 16).
+Finally, in Line 17, after the $\delta$-cliques are removed, CPGR updates $\hat{k}$ for the next iterations.
+Graph restructuring (Lines 18-23).
+When the while loop (Lines 10-17) terminates, CPGR restructures the graph by adding a new vertex set $Z$ to the bipartite graph.
+Each vertex in the left and right partitions of clique $\mathscr{C}_{q}$ is connected to a new vertex $z_{q}$, reducing the number of edges from $|U_{q}|\times|W_{q}|$ to $|U_{q}|+|W_{q}|$ (Lines 18-23).
+In Line 22, $\mathcal{E}_{\mathcal{C}}$ is updated with edges from the extracted $\delta$-cliques.
+Remaining edges in the bipartite graph $G(U,W,E)$, not part of any bipartite $\delta$-cliques, are connected directly in the tripartite graph (Line 23).
+In Line 24 the algorithm returns the restructured graph $G^{*}(U,Z,W,E^{*})$ having a reduced number of edges.
+In Appendix A.2 we provide an example showing the execution of CPGR.
+3.2 Clique Stripping Algorithm (CSA)
+The Clique Stripping Algorithm (CSA), given in Algorithm 2, extracts $\delta$-cliques from the given bipartite graph $G(U,W,E)$.
+CSA initially selects the vertices for the right partition $K_{c}$ of the clique $C_{c}$ and then the common neighbors for the set $K_{c}$.
+The input of CSA consists of the updated index $q$ of the clique, the size $\hat{k}$ of the right partition which guarantees a $\delta$-clique in $G$, the adjacency matrix $A$, the number of vertices $n$, and the degree $d_{w_{j}}=|N(w_{j})|$ .
 Algorithm 2 CSA: Clique Stripping Algorithm
-1:$q$: Index of bipartite clique;   $\hat{k}$: Size of the right bipartition in bipartite clique;
-2:      $A$: Adjacency matrix;           $n$: Number of vertices in left/right bipartition of $G$;
-3:      $d_{w}$: Degree of vertices in bipartition $W$.
-4:$\mathcal{K}\leftarrow\emptyset$
-5:Sort $d_{w}$ in non-increasing order.
-6:Let $d_{w_{\pi(1)}},d_{w_{\pi(2)}},\ldots,d_{w_{\pi(n)}}$ be the order.
-7:$j\leftarrow 1$
-8:while $d_{w_{\pi(j)}}\geq d_{w_{\pi(\hat{k})}}$ do
-9:    $\mathcal{K}\leftarrow\mathcal{K}\cup\{w_{\pi(j)}\}$
-10:    $j\leftarrow j+1$
-11:$\gamma\leftarrow\big{\lfloor}{\frac{|\mathcal{K}|}{\hat{k}}}\big{\rfloor}$
-12:for  $c=q+1,\ldots,q+\gamma$ do
-13:    $K_{c}\leftarrow\{w_{\pi_{(j)}}\in\mathcal{K}\ |\ (c-(q+1))\cdot\hat{k}<j\leq(c-q)\cdot\hat{k}\}$
-14:    $U_{K_{c}}\leftarrow\{u_{i}\in U|K_{c}\subseteq N(u_{i})\}$
-15:    $C_{c}\leftarrow\{(U_{K_{c}},K_{c})\}$
-16:    Update $A$, by removing edges $(U_{K_{c}}\times K_{c})\in C_{c}$
-17:    Update $d_{w}$ by subtracting $|U_{K_{c}}|$ from each $d_{w_{\pi(j)}}$, where $w_{\pi(j)}\in K_{c}$
-18:$\mathscr{C}\leftarrow\{C_{q+1},\ldots,C_{q+\gamma}\}$
-19:Output Set of cliques $\mathscr{C},\hat{A},\hat{d_{w}}$
+1:$q$: Index of bipartite clique;
+2:  $\hat{k}$: Size of the right bipartition in bipartite clique;
+3:  AA: Adjacency matrix;
+4:  $n$: Number of vertices in left/right bipartition of $G$;
+5:  $d_{w}$: Degree of vertices in bipartition $W$.
+[…]
+7:Sort $d_{w}$ in non-increasing order. Let $d_{w_{\pi(1)}},d_{w_{\pi(2)}},\ldots,d_{w_{\pi(n)}}$ be the order.
+[…]
+9:while $d_{w_{\pi(j)}}\geq d_{w_{\pi(\hat{k})}}$ do
+10:  $\mathcal{K}\leftarrow\mathcal{K}\cup\{w_{\pi(j)}\}$
+11:  $j\leftarrow j+1$
+[…]
+13:for $c=q+1,\ldots,q+\gamma$ do
+14:  $K_{c}\leftarrow\{w_{\pi_{(j)}}\in\mathcal{K}\ |\ (c-(q+1))\cdot\hat{k}<j\leq(c-q)\cdot\hat{k}\}$
+15:  $U_{K_{c}}\leftarrow\{u_{i}\in U|K_{c}\subseteq N(u_{i})\}$
+16:  $C_{c}\leftarrow\{(U_{K_{c}},K_{c})\}$
+17:  Update AA, by removing edges $(U_{K_{c}}\times K_{c})\in C_{c}$
+18:  Update $d_{w}$ by subtracting $|U_{K_{c}}|$ from each $d_{w_{\pi(j)}}$, where $w_{\pi(j)}\in K_{c}$
+[…]
+20:Output ($\mathscr{C},\hat{A},\hat{d_{w}}$)
 Initialization (Lines 1-3).
-CSA initializes the set of selected vertices for clique extraction, $\mathcal{K}$ to the empty set.
+CSA initializes the set $\mathcal{K}$ of selected vertices for clique extraction to the empty set.
 In Line 2, it sorts $d_{w}$ in non-increasing order of the degrees of vertices $w_{j}\in W$, i.e., $d_{w_{\pi(1)}}\geq d_{w_{\pi(2)}}\geq\ldots\geq d_{w_{\pi(n)}}$.
-Then, in Line 3 it initializes the index $j$ in the list of sorted vertices to 1.
-Vertex selection (Lines 4-7).
-CSA selects the vertices of a clique to be extracted, based on their degrees.
-Selecting a vertex with higher degree results in a larger set of neighbours that can be part of the selected clique.
-Therefore, CSA selects the vertices whose degrees are greater than or equal to the $\hat{k}$-th largest degree in the sorted $d_{w}$, i.e., $\mathcal{K}=\{w_{\pi(j)}\ |\ d_{w_{\pi(j)}}\geq d_{w_{\pi(\hat{k})}}\}$.
-CSA adds the selected vertices to $\mathcal{K}$ and increments the iterator $j$ (Lines 5-6).
-Thus, the while loop in Lines 4- 6 is executed until all vertices with degrees greater than $d_{w_{\pi(\hat{k})}}$ are chosen.
-In Line 7, CSA calculates $\gamma$ to partition the set $\mathcal{K}$ such that the size of a partition is less than or equal to $\hat{k}$.
-Therefore, every time CSA is executed it finds $\gamma$ cliques.
+In Line 3 it initializes the sorted vertices list index $j$ to 1.
+Vertex selection (Lines 4–7).
+CSA selects vertices for clique extraction based on their degrees, since vertices with higher degree are more likely to share a larger set of common neighbors and thus participate in dense bipartite cliques.
+Rather than selecting exactly $\hat{k}$ vertices, CSA constructs a candidate set $\mathcal{K}=\{w_{\pi(j)}\mid d_{w_{\pi(j)}}\geq d_{w_{\pi(\hat{k})}}\}$, which contains all vertices whose degree is at least the $\hat{k}-$th largest degree in the sorted sequence $d_{w}$.
+As a result, $|\mathcal{K}|\geq\hat{k}$ in general.
+CSA adds these vertices to $\mathcal{K}$ by iterating through the sorted order and incrementing the index $j$ (Lines 5–6).
+Thus, the while-loop in Lines 4–6 ensures that all vertices tied at the $\hat{k}-$th degree threshold are included, avoiding arbitrary tie-breaking and forming a maximal candidate set for the right partition.
+This threshold-based selection increases the likelihood of identifying a larger set of common neighbors in the left partition, thereby improving the quality of the extracted $\delta$-cliques.
+In Line 7, CSA computes $\gamma=\lfloor|\mathcal{K}|/\hat{k}\rfloor$, which determines how many $\delta$-cliques can be extracted.
+The set $\mathcal{K}$ is then partitioned into $\gamma$ subsets of size $\hat{k}$ (Line 9), allowing CSA to extract multiple $\delta$-cliques in a single execution.
 Clique extraction (Lines 8-13).
@@
 After forming the clique $C_{c}$, in Line 12, the algorithm updates the adjacency matrix by removing the edges in the clique $C_{c}$ and in Line 13, updates the degrees of the vertices that are part of the clique by removing the size of left partition from the degree of each vertex $w_{\pi(j)}\in K_{c}$, i.e., $d_{w_{\pi(j)}}=d_{w_{\pi(j)}}-|U_{K_{c}}|,\ \forall\ w_{\pi(j)}\in K_{c},\ c=q+1,\ldots q+\gamma$.
-Finally, in Line 14, CSA forms $\mathscr{C}$, the set of all the cliques extracted in the current execution, and in Line 15, it returns the set of cliques $\mathscr{C}$, the updated adjacency matrix $\hat{A}$, and the updated degrees of vertices $\hat{d}_{w}$ to CPGC.
-Time complexity of CSA.
-CSA takes $O(n\log n)$ to sort the degrees of vertices $d_{w}$, in Line 2.
-The while loop in Lines 4-6, takes at most $O(n)$ to select vertices for the $\delta$-cliques.
-The for loop (Lines 8-13) executes $O(\gamma)$ times, where $1\leq\gamma\leq\frac{n}{\hat{k}}$.
-In the for loop, Line 9 and 10 takes $O(\hat{k})$ and $O(n\hat{k})$ to find the right and left partition of the $c$-th $\delta$-clique, respectively.
-CSA takes $O(n\hat{k})$ to remove the edges in the $c$-th $\delta$-clique, in Line 12.
-It takes $O(\hat{k})$ to update the degrees $d_{w}\in K_{c}$, in Line 13.
-Therefore, the total running time of CSA is dominated by Lines 10 and 12 and thus CSA takes $O(n\gamma\hat{k})$ time.
-Time complexity of CPGC.
-CPGC takes $O(n^{2})$ to calculate the degrees $d_{w}$ of vertices in $W$ in Lines 2-4.
-The running time of the while loop (Lines 6-13) is dominated by the running time of the function CSA in Line 7 which is given by Algorithm 2.
-CSA takes $O(n\gamma\hat{k})$ to remove $\gamma\hat{k}n^{1-\delta}$ edges.
-Therefore, on average it takes $O(n^{\delta})$ time to remove one edge from the given graph.
-Thus, the while loop in CPGC takes $O(mn^{\delta})$ time to extract $\delta$-cliques.
-In the end, to compress the graph, CPGC in Lines 14-15, takes linear time.
-Therefore, the total running time of CPGC is $O(mn^{\delta})$ which is dominated by the while loop in Lines 6-13.
-4 Properties of CPGC
-Theorem 4.1
-The compressed graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGC preserves the path information of the original graph $G(U,W,E)$.
+Finally, in Line 14, CSA forms $\mathscr{C}$, the set of all the cliques extracted in the current execution, and in Line 15, it returns the set of cliques $\mathscr{C}$, the updated adjacency matrix $\hat{A}$, and the updated degrees of vertices $\hat{d}_{w}$ to CPGR.
+[…]
+[…]
+4 Properties of CPGR
+The transformation performed by CPGR preserves vertex-to-vertex reachability and the connected components of the original graph.
+Consequently, the existence of paths between any pair of vertices in $G$ is maintained in the restructured graph $G^{*}$.
+However, CPGR does not necessarily preserve local structural properties such as vertex degrees, neighborhood structure, or exact shortest-path distances, since edges belonging to $\delta$-cliques are replaced by paths of length two.
+We formalize this preservation of path information in the following theorem.
+Theorem 1.
+The restructured graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGR preserves the path information (i.e., vertex-to-vertex reachability) of the original graph $G(U,W,E)$.
 Proof.
-    Provided in Appendix A.3.
-In the following, we determine a bound on the number of edges in the compressed graph obtained by CPGC.
-This bound tells us how good the compression achieved by CPGC is.
+Let’s assume that CPGR extracts only one $\delta$-clique $C_{q}$ from the given graph $G$.
+In the $\delta$-clique $C_{q}$, the right partition $K_{q}\subseteq W$ is formed by selecting $\hat{k}$ vertices from $W$ and the left partition $U_{q}\subseteq U$ is formed with the common neighbors of $K_{q}$, forming a complete bipartite graph with edge set $E_{q}$.
+The restructured graph $G^{*}$ is formed with the same left and right partitions $U$ and $W$ of the graph $G$, and a third partition $Z$ which is a set of additional vertices associated with each of the $\delta$-cliques that CPGR extracts.
+Our main concern is with the edges $(u_{i},w_{j})\in E^{*}$ in the restructured graph $G^{*}$.
+$E^{*}$ contains the edges that replace the $\delta$-clique $C_{q}$ by adding the additional vertex $z_{q}\in Z$, and the set of edges $\hat{E}=E\setminus E_{q}$ which were not part of the $\delta$-clique $C_{q}$.
+It can be easily seen that $\hat{E}\subset E$, that is, the edges in $\hat{E}$ are edges in $G$, and the remaining edges $E^{*}\setminus\hat{E}$ in $E^{*}$ are the edges that replace $C_{q}$ in $G^{*}$.
+Each edge $(u_{i},w_{j})\in E_{q}$, is replaced by two edges, $(u_{i},z_{q})$ and $(z_{q},w_{j})$, where $u_{i}\in U$, $w_{j}\in W$, and $z_{q}\in Z$.
+Therefore, for each edge $(u_{i},w_{j})\in E_{q}$, there exists a path from $u_{i}$ to $w_{j}$ composed of two edges, $(u_{i},z_{q})$ and $(z_{q},w_{j})$, that passes through the additional vertex $z_{q}\in Z$, thus preserving the path information.
+This holds true for all the $\delta$-cliques extracted by CPGR.
+∎
+In the following, we determine an upper bound on the number of edges in the restructured graph obtained by CPGR.
 First, we state a theorem from [14] that guarantees the existence of a $\delta$-clique in a bipartite graph.
-This theorem will be used in the proofs of CPGC’s compression properties.
-Theorem 4.2
+This theorem is necessary in the proof of the bound.
+Theorem 2.
 […]
-Next, we provide a bound on the minimum number of edges $\hat{m}$ of $G$ required to obtain a compression of $G$ into $G^{*}$ by extracting $\delta$-cliques.
-Lemma 4.1
-Given a graph $G(U,W,E)$, where $n=|U|=|W|$, $m=|E|$, and a constant $\delta$, $10<\delta<$ 1, if $m<2n^{2-\frac{\delta}{2}}$ then extracting $\delta$-cliques and replacing them with tripartite graphs as done in CPGC does not lead to a compression of $G$.
+Lemma 1.
+Given a graph $G(U,W,E)$, where $n=|U|=|W|$, $m=|E|$, and a constant $\delta$, $0<\delta<1$, if $m<2n^{2-\frac{\delta}{2}}$ then extracting $\delta$-cliques and replacing them with tripartite graphs as done in CPGR does not lead to a reduction in the number of edges of $G$.
 Proof.
-    Provided in Appendix A.3.
-Theorem 4.3
-Let $G(U,W,E)$ be any bipartite graph with $|U|=|W|=n$ and $|E|=m>2n^{2-\frac{\delta}{2}}$, where $\delta$ is a constant such that $10<\delta<$ 1.
-Then, the number of edges in the compressed graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGC is $|E^{*}|=O\left(\frac{m}{k(n,m,\delta)}\right)$.
+In the $\delta$-clique based graph restructuring performed by CPGR, the restructured graph $G^{*}(U,W,Z,E^{*})$ is obtained from $G$ by replacing $|E_{q}|=|U_{q}\times K_{q}|$ edges in a $\delta$-clique $C_{q}$ with $|E^{*}|=|U_{q}|+|K_{q}|$ edges in $G^{*}$ and adding an additional vertex $z_{q}\in Z$ for each extracted $\delta$-clique $C_{q}$.
+The size of the right partition of $C_{q}$ is $|K_{q}|={k}=k(n,m,\delta)=\Big\lfloor{\frac{\delta\log n}{\log(2n^{2}/m)}}\Big\rfloor$.
+If $m<2n^{2-\frac{\delta}{2}}$ then the size of the right partition $|K_{q}|$ is $k=1$.
+Therefore, the number of edges in the $\delta$-clique is $|E_{q}|=|U_{q}|\cdot 1=|U_{q}|$.
+Those edges are replaced by $|E^{*}|=|U_{q}|+\hat{k}=|U_{q}|+1$ edges in $G^{*}$.
+Thus, replacing the $\delta$-clique in $G$ with $m<2n^{2-\frac{\delta}{2}}$, the number of edges in $G^{*}$ would actually increase by 1, which does not lead to a restructuring of $G$.
+∎
+Theorem 3.
+Let $G(U,W,E)$ be any bipartite graph with $|U|=|W|=n$ and $|E|=m>2n^{2-\frac{\delta}{2}}$, where $\delta$ is a constant such that $0<\delta<1$.
+Then, the number of edges in the restructured graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGR is $|E^{*}|=O\left(\frac{m}{k(n,m,\delta)}\right)$.
 Proof.
-    We follow the basic idea of the proof from Theorem 2.4 in [9] and extend it to apply to the CPGC algorithm.
-We assume that initially the compressed graph $G^{*}$ has 0 edges, i.e., $|E^{*}|=0$, and that edges are added to $E^{*}$ as the algorithm progress.
-To estimate the number of edges in $E^{*}$, we divide the iterations of CPGC into stages, where each stage $i$ consists of extracting one or more $\delta$-cliques with a fixed $\hat{k}_{i}$.
+We follow the basic idea of the proof from Theorem 2.4 in [14] and extend it to apply to the CPGR algorithm.
+We assume that initially the restructured graph $G^{*}$ has 0 edges, i.e., $|E^{*}|=0$, and that edges are added to $E^{*}$ as the algorithm progress.
+To estimate the number of edges in $E^{*}$, we divide the iterations of CPGR into stages, where each stage $i$ consists of extracting one or more $\delta$-cliques with a fixed $\hat{k}_{i}$.
 Therefore, the $i^{th}$ stage includes all $\delta$-cliques extracted after the number of edges in $G$ becomes less than $m/2^{i-1}$ for the first time, and before the number of edges in $G$ becomes less than $m/2^{i}$ for the first time.
@@
 The total number of edges removed from $G$ in stage $i$ cannot be greater than $2m/2^{i}$, therefore, the number of edges added to $E^{*}$ during stage $i$ is less than or equal to $\rho\frac{2m}{2^{i}}=\left(\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\right)\frac{2m}{2^{i}}$.
-    CPGC terminates extracting cliques when the number of remaining edges in $G$ is less than $2n^{2-\frac{\delta}{2}}$.
-This is to eliminate the extraction of trivial cliques, as shown in Lemma 4.1.
-Next, we determine an upper bound on the total number of edges $m^{*}$ added by CPGC to $G^{*}$ before it terminates extracting $\delta$-cliques (i.e., when $m<2n^{2-\frac{\delta}{2}}$).
+CPGR terminates extracting cliques when the number of remaining edges in $G$ is less than $2n^{2-\frac{\delta}{2}}$.
+This is to eliminate the extraction of trivial cliques, as shown in Lemma 1.
+Next, we determine an upper bound on the total number of edges $m^{*}$ added by CPGR to $G^{*}$ before it terminates extracting $\delta$-cliques (i.e., when $m<2n^{2-\frac{\delta}{2}}$).
 Therefore,
-    (4.1) $\begin{split}m^{*}\leq\sum_{i=1}^{\big{\lceil}\log\left(\frac{m}{2n^{2-\delta/2}}\right)\big{\rceil}}\bigg{(}\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\bigg{)}\frac{2m}{2^{i}}\\ \leq\sum_{i=1}^{\infty}\bigg{(}\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\bigg{)}\frac{2m}{2^{i}}.\end{split}$
+$\begin{split}m^{*}\leq\sum_{i=1}^{\big\lceil\log\left(\frac{m}{2n^{2-\delta/2}}\right)\big\rceil}\bigg(\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\bigg)\frac{2m}{2^{i}}\leq\sum_{i=1}^{\infty}\bigg(\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\bigg)\frac{2m}{2^{i}}.\end{split}$ (1)
 Since $1/(2k_{i})\leq 1/k+i/(\delta\log n)$, we obtain,
-(4.2) $\begin{split}m^{*}\leq\frac{4m}{k}\sum_{i=1}^{\infty}\frac{1}{2^{i}}+\frac{4m}{\delta\log n}\sum_{i=1}^{\infty}\frac{i}{2^{i}}+\frac{2m}{n^{1-\delta}}\sum_{i=1}^{\infty}\frac{1}{2^{i}}\\ \leq 2m\left(\frac{2}{k}+\frac{4}{\delta\log n}+\frac{1}{n^{1-\delta}}\right).\end{split}$
+$\begin{split}m^{*}\leq\frac{4m}{k}\sum_{i=1}^{\infty}\frac{1}{2^{i}}+\frac{4m}{\delta\log n}\sum_{i=1}^{\infty}\frac{i}{2^{i}}+\frac{2m}{n^{1-\delta}}\sum_{i=1}^{\infty}\frac{1}{2^{i}}\leq 2m\left(\frac{2}{k}+\frac{4}{\delta\log n}+\frac{1}{n^{1-\delta}}\right).\end{split}$ (2)
 Thus, $m^{*}=O\left(\frac{m}{k}\right)$.
-After CPGC finishes extracting cliques, there are $2n^{2-\delta/2}$ edges remaining in $G$.
+After CPGR finishes extracting cliques, there are $2n^{2-\delta/2}$ edges remaining in $G$.
 Those remaining edges are trivial cliques (i.e., single edges) that are added to $G^{*}$.
 The number of remaining edges $2n^{2-\delta/2}$ is in $O\left(\frac{m}{k}\right)$.
-Therefore, the number of edges in the compressed graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGC is $|E^{*}|=O\left(\frac{m}{k}\right)$.
-Extension to non-bipartite graphs. In Appendix A.4 we describe how CPGC can be extended to compress non-bipartite graphs.
+Therefore, the number of edges in the restructured graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGR is $|E^{*}|=O\left(\frac{m}{k}\right)$.
+∎
+Extension to non-bipartite graphs.
+CPGR can be extended to restructure non-bipartite graphs using a similar technique to that described in [14].
+Consider a directed graph $H(V,E)$ with $n$ vertices and $m$ edges.
+Initially, we transform $H$ into a bipartite graph $G\left(L,R,E^{\prime}\right)$, where each vertex $v\in V$ is duplicated into $v_{L}\in L$ and $v_{R}\in R$.
+For each directed edge $(u,v)$ in $E$, we create a directed edge $\left(u_{L},v_{R}\right)$ in $E^{\prime}$, with the direction from $L$ to $R$.
+Furthermore, for each vertex $v\in V$ we add a directed edge from $v_{R}$ to $v_{L}$.
+The proposed graph restructuring algorithm is then applied to the bipartite graph $G$, which includes only the edges from $L$ to $R$.
+Once the restructured graph is computed we add to it the $n$ directed edges from $u_{R}$ to $u_{L}$, corresponding to each vertex $u\in V$.
+Since the path information of the original graph is preserved by this transformation, it allows different graph algorithms to work on the restructured graph as well.
+Undirected general graphs can be transformed first into directed graphs by simply replacing each undirected edge $(u,v)$ in the original graph by two directed edges $(u,v)$ and $(v,u)$, and then applying the technique presented above to transform the directed graph into a bipartite graph.
+The time complexity and the compression ratio of the proposed algorithms are the same as in the case of the bipartite graphs.
+The time complexity and the reduced number of edges of CPGR are the same as in the case of the bipartite graphs.
+A similar approach for converting a general graph to a bipartite graph and then restructuring the obtained bipartite graph was employed in [3, 10].
 5 Experimental Results
-We investigated the performance of CPGC in terms of both running time and compression ratio, and compared it with the FM algorithm proposed by Feder and Motwani [9].
-We could not locate any existing implementation of the FM algorithm.
-Thus, we independently implemented it in C to the best of our ability.
-However, we encountered a limitation with the FM algorithm regarding large graphs, where the calculations for selecting vertices may exceed the machine representation limits.
-Consequently, we could only obtain results for bipartite graphs with up to $n=128$ vertices in each bipartition and 16 thousand edges (i.e., small graphs).
-To ensure a fair comparison between algorithms, we excluded the time spent on extracting trivial cliques ($\hat{k}=1$) when measuring the runtime of the FM algorithm.
-CPGC addresses this by excluding trivial cliques considering $\hat{k}>1$, as indicated in Line 6 of Algorithm 1.
-Furthermore, we investigated the performance of our algorithm, CPGC, on large bipartite graphs with up to 32 thousand vertices in each bipartition and approximately 1.05 billion edges.
-Both CPGC and FM algorithms were implemented in C and experiments were conducted on a Linux system with an AMD EPYC-74F3 processor, 1 CPU core, 3.2GHz, and 128 GB of memory.
-We used the GCC compiler (version 8.5.0) for compiling and executing the C code.
-Figure 3: CPGC: Average compression ratio (top row: (a), (b), and (c)) and running time (bottom row: (d), (e), and (f)) for graphs with 3.36 million to 1.05 billion edges.
-For our experiments, we generated bipartite graphs with various densities using the $G(n,p)$ random graph model adapted to bipartite graphs.
-In this model, an edge is included in the bipartite graph with probability $p$, independently from every other edge.
-We implemented a Python program to generate instances of bipartite graphs with the number of vertices $|U|=|W|=n$, where $n$ ranges from 32 to 32,768, and densities of 0.8, 0.85, 0.90, 0.95, and 0.98.
-This generated graphs with up to 1.05 billion edges, considering $p$ as a measure of the density of the generated bipartite graphs.
-For each bipartite graph with a given density, we created 10 different instances.
-We then ran each instance for six different values of delta ($\delta$: 0.5, 0.6, 0.7, 0.8, 0.9, 1), presenting the average values and standard deviations of the running time and compression ratio for these 10 instances.
-5.1 Results for Large Graphs.
-We present the results obtained from testing CPGC on large bipartite graphs, where each bipartition consists of $n=2^{i}$ vertices, with $i$ ranging from 11 to 15, and a number of edges $m$ ranging from approximately 3.36 million to 1.05 billion.
-Figures 3a-3c show the average compression ratio, while Figures 3d-3f show the average running time for bipartite graphs.
-Average compression ratio.
-The compression ratio for CPGC is calculated as $\frac{m}{m^{*}}$, where $m$ and $m^{*}$ represent the number of edges in the given and compressed graphs, respectively.
-We observe that, for the same $n$ and $\delta$, the compression ratio increases with density.
-For instance, with $n=2^{15}$ and $\delta=0.5$, at a density of $0.80$ with $m=214.75$ million, the compression ratio is $1.95$.
-In contrast, at a density of $0.98$, with $m=1.05$ billion, the compression ratio increases to $2.44$.
-This is because the probability of finding common neighbours is correlated with the density, that is, increasing the density increases the probability of finding a large set of common neighbours of the right partition of the $\delta$-clique.
-Finding more common neighbours results in CPGC extracting more edges, and thus increasing the compression ratio.
-However, when the density is kept constant, the compression ratio is not monotonic and depends on $\delta$ and $n$.
-For instance in Figure 3a, when $n=2^{13}$ and $m=53.69$ million for density $0.80$, the compression ratio initially increases from 1.85, for $\delta=0.5$, to 2, for $\delta=0.6$, achieving the maximum compression.
-Yet, further increasing $\delta$ to 1 leads to a reduction in the compression ratio to 1.36.
-This is because increasing $\delta$ increases the number of vertices in the right partition $K_{q}$ of a $\delta$-clique, as $|K_{q}|=\hat{k}$ and $\hat{k}\propto\delta$.
-Consequently, this decreases the probability of finding the common neighbors, and therefore for lower density, increasing $\delta$ results in a decrease in the compression ratio.
-Conversely, as the density increases, the probability of discovering common neighbors also increases.
-Hence, at a density of 0.98, an increase in $\delta$ correlates with an increase in the compression ratio for all $n$ and $m$ values (Figure 3c).
-It is important to note that the compression ratio depends on $n$, $m$, and $\delta$.
-Certain combinations of $n$, $m$, and $\delta$ lead to the highest compression ratio.
-We leave the investigation of finding the optimal combination of these parameters for future work.
-Figure 4: CPGC vs. FM: Compression ratio relative to FM (top row: (a), (b) and (c)) and speedup relative to FM (bottom row: (d), (e) and (f)) for graphs with 819 to 16 thousand edges.
-Average running time.
-We observe that for the same $\delta$ and density, the running time increases as the number of vertices increases which is expected (CPGC running time is $O(mn^{\delta})$).
-We also observe that for the same $n$ and density, the running time decreases as $\delta$ increases.
-For instance, with $m=214.75$ million for density = 0.80 and $\delta=0.5$, the running time is 176.27 seconds, whereas for $\delta=1$ it reduces to 117.8 seconds.
-This is because of the relation between $\delta$ and $\hat{k}$ described earlier, CPGC cannot find a large set of common neighbours which results in a lower compression ratio for lower density.
-Also, when CPGC does not find an edge with $u_{i}\in U$ for any vertex $w_{j}\in K_{q}$, then it skips to the next vertex in $U$ to check for an edge with $w_{j}\in K_{q}$ thus, decreasing the running time.
-Therefore, we do not observe an increase in the running time as $\delta$ increases.
-However, for density = 0.98, the running time decreases as $\delta$ increases, even when the compression ratio increases.
-For example, with $m=1.05$ billion for density = 0.80 and $\delta=0.5$, the running time is 164.83 seconds, whereas for $\delta=1$ the running time is 150.97 seconds.
-This is because for higher density and large $\delta$, large $\delta$-cliques are extracted which decreases the total number of iterations performed by CPGC and subsequently reduces the running time.
-Figure 5: Dinitz($G$) vs.
-Dinitz($G^{*}$): Average running time of Dinitz’s algorithm on the original bipartite graph $G$ (labeled Dinitz($G$)) and on the compressed graph $G^{*}$ (labeled Dinitz($G^{*}$)), for a large graph with approximately 32,000 vertices in each bipartition and 214.75 million to 1.05 billion edges corresponding to densities = $0.80,0.90,$ and $0.98$ and different $\delta$.
-5.2 Results for Small Graphs.
-In this section, we compare the performance of CPGC against FM on small graphs, particularly, on graphs with $n=2^{i}$ vertices in each bipartition, where $i$ ranges from 5 to 7 and the number of edges ranging from 819 to 16,056.
+We evaluated the running time and the reduction in the number of edges, $R$ of CPGR by comparing it with the FM algorithm [14].
+Let $m=|E|$ be the number of edges in the given graph $G$ and $m^{*}=|E^{*}|$ be the number of edges in the restructured graph $G^{*}$ and $m^{*}\leq m$, then we define the reduction in the number of edge as $R=m-m^{*}$.
+Due to the lack of an existing implementation, we implemented FM in C.
+However, its vertex selection calculations encountered machine representation limits on large graphs, restricting our comparative analysis to small bipartite graphs (up to $n=128$ vertices per partition, $\sim$16k edges).
+To ensure a fair runtime comparison, we excluded trivial clique extraction ($\hat{k}=1$) for both algorithms (CPGR inherently does this as per Algorithm 1, Line 10).
+We further assessed CPGR on large bipartite graphs (up to 1.05 billion edges) and real-world datasets from the SuiteSparse Matrix Collection [9].
+Additionally, we tested the utility of our restructured graph by applying the Dinitz algorithm for maximum bipartite matching to large bipartite graphs and the APSP algorithm on real-world datasets.
+All algorithms, were implemented in C and tested on a Linux system (AMD EPYC-74F3, 3.2GHz, 1 core, 128 GB RAM) using GCC 8.5.0.
+The implementation is available at GitHub [17].
+Our experiments involved synthetic bipartite graphs generated using adapted $G(n,p)$ random graph model with $|U|=|W|=n$, $n\in[32,32768]$, densities $p\in[0.8,0.98]$, up to 1.05 billion edges and general graphs with standard $G(n,p)$ with $n=32000$, densities $p\in[0.4,0.7]$, $\sim$409 to 716 million edges, with 10 instances per configuration.
+We also used real-world sparse graphs from SuiteSparse [9].
+For each instance, we ran CPGR with six $\delta$ values [0.5 to 1] and show average results and standard deviations for runtime and reduction in number of edges $R$, as well as the performance of the application algorithms on the restructured graphs.
+5.1 Results for Small Bipartite Graphs
+Figure 3: CPGR vs. FM: Reduction in number of edges relative to FM (top row: (a), (b) and (c)) and speedup relative to FM (bottom row: (d), (e) and (f)) for graphs with 819 to 16 thousand edges.
+In this section, we compare the performance of CPGR against FM on small bipartite graphs, particularly, with $n=2^{i}$ vertices in each bipartition, where $i$ ranges from 5 to 7 and the number of edges ranging from 819 to 16,056.
 However, due to the limitations of FM as previously mentioned, our comparison is constrained to $n=2^{7}$ and $m=16,056$.
-We compare the performance of the two algorithms using two metrics, relative compression ratio, and relative speedup.
-Compression ratio relative to FM.
-We define the compression ratio relative to FM, as the ratio of compression ratio achieved by CPGC and FM.
-Figures 4a-4c show the relative compression ratio relative to FM achieved by CPGC in the case of small graphs, for $\delta=0.5,0.6,0.7,0.8,0.9,1$, and densities = 0.80, 0.90, 0.98.
-In all cases, the relative compression ratio ranges from $1$ to $1.27$, highlighting that the compression ratio achieved by CPGC is at least as that of FM.
-CPGC achieves a greater compression ratio than FM because it extracts multiple $\delta$-cliques before $\hat{k}$ decreases as shown in Figure 2(b), resulting in removing more edges at higher values of $\hat{k}$ (Section 3.2).
-We observe that CPGC achieves a compression ratio equal to that of FM for a small graph with $n=32$, $m=819$, density = 0.80, and $\delta=0.5$.
-This is because a small graph does not require many iterations for extracting $\delta$-cliques, thereby not giving CPGC the opportunity to extract more cliques over the FM algorithm.
-However, for the same $n$ and $\delta$ when the graph density increases the compression ratio achieved by CPGC also increases.
-For example, with $n=32$, $m=921$, and $1,003$, the relative compression ratio is 1.03 and 1.25, respectively.
-As explained before, certain combinations of $n$, $m$, and $\delta$ lead to the highest compression ratio.
-For instance, when $n=128$, $m=16,056$ for a density of 0.98, we observe that the relative compression ratio initially increases from 1.1 to 1.14 with increasing $\delta$ from 0.5 to 0.7.
-However, further increasing $\delta$ to 1 results in a decrease in the relative compression ratio to 1.1.
+We compare the performance of the two algorithms using two metrics, speedup, and $R$ relative to FM.
+Reduction in number of edges $R$ relative to FM.
+We define the edge reduction ratio relative to FM as the ratio of the $R$ achieved by CPGR to that of FM.
+Figures 3a-3c illustrate the relative $R$ of CPGR compared to FM for small bipartite graphs across different $\delta$ values (0.5 to 1) and densities (0.80, 0.90, 0.98).
+In all these cases, the relative $R$ ranges from 1 to 1.27, indicating that CPGR achieves at least the same, and often a greater reduction than FM.
+This improved performance of CPGR stems from its ability to extract multiple $\delta$-cliques before $\hat{k}$ decreases (as shown in Figure 2b), allowing for the removal of more edges at higher $\hat{k}$ values (Section 3.2).
+Notably, for a graph with $n=32$, $m=819$, a density of 0.80, and $\delta$=0.5, CPGR achieves an $R$ equivalent to FM.
+This is because the small size of the graph limits the number of iterations required for $\delta$-clique extraction, thus not providing CPGR with a significant advantage over FM.
+However, for the same $n=32$ and $\delta$, increasing the graph density leads to a higher $R$ for CPGR.
+For instance, when $m$ increases from 921 to 1,003, the relative $R$ rises from 1.03 to 1.25.
+In contrast, for a graph $n=128$, $m=16056$, density = 0.98, we observe a non-monotonic trend in the relative $R$ as $\delta$ increases.
+Initially, the ratio increases from 1.1 to 1.14 as $\delta$ goes from 0.5 to 0.7.
+However, further increasing $\delta$ to 1 results in a decrease to 1.1.
+This reduction is to the fact that a larger $\delta$ increases the size of the right partition $K_{q}$ of a $\delta$-clique (since $|K_{q}|=\hat{k}$ and $\hat{k}\propto\delta$), consequently decreasing the likelihood of finding common neighbors.
+It is important to recognize that the optimal $R$ is influenced by the interplay between $n,m,$ and $\delta$.
+Identifying the specific combinations of these parameters that yield the highest reduction in the number of edges $R$ remains a subject for future investigation.
 Speedup relative to FM.
-We define the speedup relative to FM as the ratio of the runing time of FM over CPGC.
-Figures 4d-f illustrate the speedup achieved by CPGC for bipartite graphs with $n$ = 32, 64, and 128 vertices in each partition, $m$ = 819 to 16 thousand for densities of 0.80, 0.90, and 0.98, and $\delta$ values ranging from 0.5 to 1.
-The running time of CPGC is consistently lower than that of FM, resulting in speedup by CPGC across all cases.
-This is primarily due to the selection of vertices in the while loop (Lines 4-6) in CSA, facilitating faster compression by extracting more than one i.e., $\gamma$, $\delta$-cliques in a single iteration, as shown in Figure 2(b).
+We define the speedup relative to FM as the ratio of the runing time of FM over that of CPGR.
+Figures 3d-f illustrate the speedup achieved by CPGR for bipartite graphs with $n$ = 32, 64, and 128 vertices in each partition, $m$ = 819 to 16 thousand for densities of 0.80, 0.90, and 0.98, and $\delta$ values ranging from 0.5 to 1.
+The running time of CPGR is consistently lower than that of FM, resulting in speedup by CPGR across all cases.
+This is primarily due to the selection of vertices in the while loop (Lines 4-6) in CSA, facilitating greater $R$ by extracting more than one (i.e., $\gamma$), $\delta$-cliques in a single iteration, as shown in Figure 2b.
 The speedup increases notably with higher $\delta$ and density.
-For example, with $n=32$ and $m=819$ for density 0.8, CPGC achieves an average speedup of 3.66 for $\delta=0.5$, increasing to 16.28 for $\delta=1$.
-Similarly, with $n=32$ and $\delta=1$, CPGC achieves an average speedup ranging from 16.28 for $m=819$ with density 0.8 to 20.97 for $m=16$ thousand with density 0.98.
+For example, with $n=32$ and $m=819$ for density 0.8, CPGR achieves an average speedup of 3.66 for $\delta=0.5$, increasing to 16.28 for $\delta=1$.
+Similarly, with $n=32$ and $\delta=1$, CPGR achieves an average speedup ranging from 16.28 for $m=819$ with density 0.8 to 20.97 for $m=16$ thousand with density 0.98.
 It is worth noting that for small graphs, certain combinations of $n$, $m$, and $\delta$ yield higher speedup values.
 For instance, with $n=128$ and $m=819$ for density 0.8, the speedup increases from 52.81 to 81.97 for $\delta=0.8$, and then slightly decreases to 60.68 for $\delta=1$.
-5.3 Speeding-up Matching Algorithms.
-Since the compressed representation of the original graph maintains all the path information, it can be used as input to other graph algorithms such as maximum cardinality matching, to speedup their execution.
-Consider Dinitz’s algorithm [6] for finding the maximum cardinality matching in a bipartite graph, whose running time is in $O(\sqrt{n}m)$.
-According to Theorem 4.3 the compression ratio of CPGC is $k=\Omega\Big{(}\frac{\delta\log n}{\log(2n^{2}/{m})}\Big{)}$, where $10<\delta<$ 1.
-This is the same as the compression ratio achieved by FM [9].
-Because the number of edges in the compressed graph is reduced by a factor of $k$, using CPGC as preprocessing for Dinitz’s algorithm leads to a total time of $O(\sqrt{n}m/k+mn^{\delta})$.
-For $\delta<1/2$, the total running time becomes $O(\sqrt{n}m\log_{n}{(\frac{n^{2}}{m}}))$.
-To the best of our knowledge, this matches the best known asymptotic bound for maximum cardinality bipartite matching algorithms (also achieved by employing FM instead of CPGC and by the bipartite matching algorithm based on push-relabel method [10]).
-Similar reasoning applies if instead of Dinitz’s algorithm we consider the Hopcroft-Karp algorithm [11].
-Figure 5 shows the average running time of Dinitz’s algorithm on the original bipartite graph $G$ with $n=32$ thousand vertices in each bipartition, and $m$ ranging from 214.75 million to 1.05 billion edges for densities = $0.80,0.90,$ and $0.98$ and different $\delta$.
-It also shows the run time of Dinitz’s algorithm executed on the compressed tripartite graph $G^{*}$, for different values of $\delta$, and densities $=0.80,0.90,$ and $0.98$.
-We observe that in all cases the Dinitz algorithm applied to the compressed graph $G^{*}$ outperforms its execution on the original graph $G$, achieving a reduction in the running time of up to 72.83%.
-Even when considering the preprocessing time required to compress the graph $G$, the running time of the Dinitz algorithm for large, dense graphs is still reduced by up to 25.09% with the compressed graph $G^{*}$.
-This reduction in the running time is directly proportional to the level of compression achieved, where larger compression ratios result in reduced running time.
-For instance, as discussed in the context of large graphs (Figure 3(a-c)), for a density of $0.98$, CPGC demonstrates an increasing compression ratio as the edge count $m$ reaches 1.05 billion with increasing values of $\delta$.
-Consequently, the run time of Dinitz’s algorithm on the compressed graph $G^{*}$ also decreases.
+5.2 Results for Large Bipartite Graphs
+We present the results obtained from testing CPGR on large bipartite graphs, where each bipartition consists of $n=2^{i}$ vertices, with $i$ ranging from 11 to 15, and a number of edges $m$ ranging from approximately 3.36 million to 1.05 billion.
+Figure 4: CPGR on Large Graphs: Average reduction in the number of edges $R^{avg}$ ((a), (b), and (c)) and average running time ((d), (e), and (f)) for graphs with 3.36 million to 1.05 billion edges.
+Average reduction in the number of edges ($R^{avg}$).
+Figures 4a-4c illustrate $R^{avg}$ for CPGR, defined as the percentage of edges removed from the original graph to obtain the restructured graph, i.e., $R^{avg}=\frac{m-m^{*}}{m}\cdot 100$.
+We observe a consistent trend: for fixed $n$ and $\delta$, $R^{avg}$ improves with increasing density.
+For example, when $m=214.75$ million edges at a density of $0.80$ and $\delta=0.5$, $R^{avg}$ is $48.72\%$, whereas $R^{avg}$ rises to $59\%$ for $m=1.05$ billion edges at a density of $0.98$.
+This correlation arises because higher density increases the likelihood of finding a large set of common neighbors for the right partition of a $\delta$-clique.
+Identifying more common neighbors allows CPGR to remove more edges, thereby improving $R^{avg}$.
+However, when the density is held constant, $R^{avg}$ exhibits a non-monotonic behavior depending on $\delta$ and $n$.
+Consider Figure 4a, for $m=53.69$ million edges at density $0.80$, $R^{avg}$ initially increases from $45.6\%$ at $\delta=0.5$ to a peak of $50\%$ at $\delta=0.6$.
+Further increasing $\delta$ to $1$, however, leads to a decrease in $R^{avg}$ to $26.5\%$.
+As explained before, this is because a larger $\delta$ increases the size of the right partition $K_{q}$ of a $\delta$-clique, which consequently lowers the possibility of finding common neighbors.
+Average running time.
+For a fixed $\delta$ and density, the running time of CPGR increases with the number of vertices, consistent with its $O(mn^{\delta})$ complexity.
+Conversely, for a fixed $n$ and density, the running time decreases as $\delta$ increases.
+For example, at density $0.80$ ($m=214.75$ million), the runtime drops from $176.27$ s at $\delta=0.5$ to $117.8$ s at $\delta=1$.
+This is because a larger $\delta$ limits the search for common neighbors, leading to lower reduction in the number of edges at lower densities and allowing CPGR to skip more vertices.
+Consequently, the expected increase in runtime with $\delta$ is not observed.
+Notably, even at a high density of $0.98$ ($m=1.05$ billion), the running time still decreases with increasing $\delta$ (from $164.83$ s at $\delta=0.5$ to $150.97$ s at $\delta=1$).
+This is attributed to the extraction of larger $\delta$-cliques at higher densities and $\delta$, which reduces the overall number of iterations in CPGR, thus lowering the runtime.
+5.3 Results for General Graphs
+Figure 5: CPGR on General Graphs: (a) average reduction in number of edges $R^{avg}$ and (b) average running time for general graphs with 409 million to 716 million edges.
+In this section, we analyze CPGR’s performance on synthetically generated general graphs with $n=32000$ vertices and number of edges $m$ ranging from approximately 409 million to 716 million as explained in Section 5.
+We observe that $R^{avg}$ and running time follow the same trend with varying density and $\delta$ as seen in the large bipartite graphs.
+Average reduction in the number of edges ($R^{avg}$).
+Figure 5a shows that for a fixed $\delta$, $R^{avg}$ improves with density.
+For instance, at $\delta=0.6$, $R^{avg}$ increases from $12.3\%$ for $m\approx 409$ million at density $0.4$ to $28.6\%$ for $m\approx 716$ million at density $0.7$, consistent with the increased likelihood of finding common neighbors.
+Similar to large bipartite graphs, $R^{avg}$ is non-monotonic with $\delta$ at a constant density.
+For $m\approx 716$ million at density $0.7$, $R^{avg}$ peaks at $33.3\%$ for $\delta=0.6$, after initially increasing from $28.6\%$ at $\delta=0.5$, before dropping to $9.1\%$ at $\delta=1$.
+Average running time.
+Running time exhibits a non-monotonic trend with increasing $\delta$ and density.
+Larger $\delta$ leads to larger right partitions, reducing the likelihood of finding common neighbors in low-density graphs, thus fewer extracted cliques with larger partitions.
+Conversely, low $\delta$ (e.g., 0.5) results in smaller right partitions, increasing the chance of finding common neighbors, leading to higher $R$ and longer runtime, as evident in Figure 5b.
+The highest runtime (59.96 s) corresponds to the highest $R^{avg}$ ($33.3\%$) at density $0.7$ and $\delta=0.6$, while the runtime reduces to 21.21 s when $\delta=1$, and $R^{avg}$ drops to $9.1\%$ at the same density.
+Table 2: Real-world Datasets (SuiteSparse Matrix Collection [9])
+Sr. No Dataset name Rows Columns Edges Density
+1 beacxc 497 506 50409 0.2
+2 beaflw 497 507 53403 0.2
+3 cari 400 1200 152800 0.3
+4 mbeacxc 496 496 49920 0.2
+5 mbeaflw 496 496 49920 0.2
+Figure 6: CPGR on Real-world graphs: (a) average reduction in the number of edges ($R^{avg}$) and (b) average running time.
+5.4 Results for Real-world Graphs
+In this section, we analyze CPGR performance on low-density (up to 0.3), symmetric and asymmetric real-world graphs shown in Table 2.
+Given that $\delta<0.8$ would yield $\hat{k}=1$ and thus no reduction in the number of edges for these sparse graphs, we tested CPGR with $\delta\in\{0.8,0.9,1\}$.
+Average reduction in the number of edges ($R^{avg}$).
+Figure 6a demonstrates that CPGR achieves $R^{avg}$ ranging from $18.7\%$ to $46.8\%$ even for the relatively low-density real-world graphs examined.
+Notably, for dataset cari, $R^{avg}$ remains consistent across all tested $\delta$ values.
+This stability is due to low density, which limits the maximum value of $\hat{k}$ to 2, thereby resulting in same the $R^{avg}$ regardless of $\delta$.
+While for other real-world graphs, $R^{avg}$ increases for $\delta$ = 1, this is because, when $\delta=1$, $\hat{k}$ larger for longer iterations of CPGR (as shown in Figure 2b), resulting in increase in $R^{avg}$.
+Average running time.
+The running time of CPGR on the real-world datasets is presented in Figure 6b.
+Similar to $R^{avg}$, the running time generally remains stable for most datasets, which is due to the same reason of low density limiting $\hat{k}$.
+However, we observe for the datasets that show increase in $R^{avg}$ at $\delta$ = 1, has increase in running time, For example, the average running time for dataset beaflw increases from 0.007 s at $\delta=0.8$ and $0.9$ to 0.014 s at $\delta=1$.
+This increase in running time is because at $\delta=1$, CPGR could extract larger $\delta$-cliques (with $\hat{k}=2$), leading to higher $R^{avg}$ and thus a longer running time.
+5.5 Speeding-up Graph Algorithms
+In this section, we show the speedup achieved by downstream graph algorithms when we use CPGR algorithm as a preprocessing step across different graph types.
+Specifically, we present the speedup of the APSP algorithm on real-world datasets and the speedup of Dinitz’s algorithm on large synthetic bipartite datasets.
+Across these algorithms and dataset types, utilizing graphs with reduced number of edges obtained by CPGR as input consistently results in smaller running times.
+Figure 7: CPGR + APSP($G^{*}$) vs.
+APSP($G$): Average running time of APSP algorithm on the original graph $G$ (labeled APSP($G$)) and on the restructured graph $G^{*}$ (labeled APSP($G^{*}$)), for large graphs with approximately 32,000 vertices and 256 million to 360 million edges corresponding to densities = $0.50,0.60,$ and $0.70$ and different $\delta$.
+Speeding-up All-Pairs-Shortest-Path Algorithms.
+We evaluate the effectiveness of CPGR for speeding up Breadth-First Search (BFS)–based All-Pairs Shortest Path (APSP) computations on unweighted general graphs.
+Figure 7 shows the running time of APSP on the original graph $G$ and on the restructured graph $G^{*}$ obtained via CPGR.
+For each density, the bars represents the average cost of CPGR, APSP($G^{*}$), and APSP($G$), measured over 10 runs.
+To validate correctness, we compare the number of connected vertices in the original and restructured graphs.
+This validation confirms that vertex-to-vertex reachability is preserved between $G$ and $G^{*}$.
+Moreover, any path discovered by BFS in the restructured graph $G^{*}$ can be mapped back to a corresponding path in the original graph $G$ by contracting the length-two paths introduced by CPGR for $\delta$-cliques.
+Therefore, while path lengths in $G^{*}$ may differ, the APSP results remain correct with respect to the original graph.
+Figure 7 shows the results on synthetic graphs with $n=32000$ vertices, varying both graph density and the parameter $\delta$.
+The results show that, across all tested densities, the combined execution time of CPGR and APSP($G^{*}$) consistently outperforms APSP($G$) on large graphs.
+Although the preprocessing cost increases with density, the reduction in APSP runtime offsets this overhead, yielding speedups of up to 1.74x.
+These results demonstrate that CPGR remains effective even for large, moderately dense graphs.
+Speeding-up Matching Algorithms.
+Figure 8: CPGR + Dinitz($G^{*}$) vs.
+Dinitz($G$): Average running time of Dinitz’s algorithm on the original bipartite graph $G$ (labeled Dinitz($G$)) and on the restructured graph $G^{*}$ (labeled Dinitz($G^{*}$)), for a large graph with approximately 32,000 vertices in each bipartition and 819.75 million to 1.05 billion edges corresponding to densities = $0.80,0.90,$ and $0.98$ and different $\delta$.
+The restructured graph retains path information, enabling its use as input for other graph algorithms like maximum cardinality matching to accelerate execution.
+For bipartite matching, we focus on Dinitz’s algorithm, which computes a maximum matching via the standard unit-capacity flow formulation.
+In this formulation, each edge has unit capacity and feasible flows correspond exactly to matching.
+Under the CPGR restructuring, each edge $(u_{i},w_{j})$ belonging to an extracted $\delta$-clique is replaced by a two-edge path $(u_{i},z_{q},w_{j})$ in the restructured graph $G^{*}$, where both edges have unit capacity.
+This transformation preserves the set of feasible flows: a flow can traverse $(u_{i},z_{q},w_{j})$ in $G^{*}$ if and only if the corresponding edge $(u_{i},w_{j})$ could be used in $G$, and matching constraints continue to be enforced by the unit-capacity edges incident to vertices in the left and right partitions.
+Consequently, Dinitz’s algorithm computes a maximum flow of the same value on $G^{*}$ as on $G$, and the resulting matching can be mapped back to a maximum matching in $G$ in linear time.
+Dinitz’s algorithm [11], with a runtime of $O(\sqrt{n}m)$ for bipartite matching, benefits from CPGR preprocessing.
+Given CPGR’s reduction in the number of edges $k=\Omega\Big(\frac{\delta\log n}{\log(2n^{2}/{m})}\Big)$ (matching FM’s), the total time becomes $O(\sqrt{n}m/k+mn^{\delta})$, which simplifies to $O(\sqrt{n}m\log_{n}{(\frac{n^{2}}{m}}))$ for $\delta<1/2$.
+Notably, this matches the asymptotic bound for maximum cardinality bipartite matching, achieved by both FM and push-relabel methods.
+Similar speedups are expected with the Hopcroft-Karp algorithm [16].
+Better asymptotic bounds (almost linear in $m$) are achieved by recent bipartite cardinality matching algorithms such as those presented in [8] ($n^{2+o(1)}$-time randomized algorithm) and [21] ($\tilde{O}(m+n^{1.5})$-time randomized algorithm).
+These algorithms are mainly of theoretical interest, providing the best bounds, and not of practical interest.
+This is due to the use of sophisticated techniques that require significant implementation effort and they might not even be feasible to implement in practice.
+Thus, for our experimental analysis we use Dinitz’s algorithm as the baseline for comparisons.
+Figure 8 compares the running time of Dinitz’s algorithm on the original bipartite graph $G$ (n=32n=32k per partition, $m=214.75$ Million to 1.05 Billion edges, densities 0.80-0.98) with the combined time of CPGR preprocessing and Dinitz on the restructured graph $G^{*}$.
+In all cases, the combined time outperforms Dinitz on $G$, achieving up to a 51.63% reduction (2.07x speedup).
+On average CPGR + Dinitz ($G^{*}$) achieves 1.69x speedup when compared to Dinitz ($G$).
+The speedup correlates with the reduction in the number of edges; for example, at density 0.98, the increasing reduction in the number of edges by CPGR with higher $m$ and $\delta$ leads to a faster Dinitz on $G^{*}$, resulting in up to a 96.07% reduction in Dinitz’s runtime compared to $G$.
 6 Conclusion and Future Work
-We proposed a new algorithm for graph compression based on partitioning the graph into bipartite cliques.
-The proposed algorithm runs in time $O(mn^{\delta})$ which improves over the $O(mn^{\delta}\log^{2}n)$ time of the Feder-Motwani (FM) algorithm (the best existing clique partitioning-based graph compression algorithm).
-Our algorithm achieves an average compression ratio at least as good as that of FM.
-Experimental results showed that our algorithm achieved a compression ratio of up to 26% greater than that obtained by FM and executes up to 105.18 times faster than FM.
-We also investigated the reduction in running time of a cardinality matching algorithm (Dinitz’s algorithm) when using the compressed graph instead of the original graph.
-The results showed that for sufficient large dense graphs, using the compressed graph obtained by CPGC as an input to the matching algorithm leads to a reduction in the running time of up to 72.83% over the matching algorithm using the original graph.
-In future work, we plan to continue investigating the graph compression problem and propose algorithms that improve over the compression ratio obtained by our proposed algorithm.
-A Appendix
-A.1 FM algorithm.
+We proposed a novel Clique Partition based Graph Restructuring (CPGR) algorithm that operates by partitioning the graph into bipartite cliques.
+The algorithm has a time complexity of $O(mn^{\delta})$.
+Experimental results demonstrate the effectiveness of CPGR, achieving a reduction in the number of edges by up to 21% greater and execution times up to 105.18 times faster than the baseline FM algorithm.
+Furthermore, we investigated the impact of using the CPGR-restructured graph as input for a cardinality matching algorithm (Dinitz’s algorithm) and an All-Pairs Shortest Path (APSP) algorithm.
+Experimental results show that for sufficiently large dense graphs, our approach leads to a reduction in the total running time (including CPGR time) of up to 51.63%, resulting in a speedup of 2.07x over the matching algorithm on the original graph.
+Similarly, for the APSP algorithm, we observed up to 34.78% reduction in total running time, achieving a speedup of up to 1.74x.
+These speedups for fundamental graph algorithms highlight the practical benefits of our graph restructuring technique, particularly for large-scale real-world applications like route planning and social network analysis where faster processing is critical for efficiency and responsiveness.
+Looking ahead, our results indicate that the edge reduction achieved by CPGR is influenced by the interaction between the graph size, the graph density, and the parameter $\delta$.
+A deeper understanding of these interactions, including the behavior observed for certain dense graphs, remains an open problem.
+Future work will focus on theoretically characterizing these interactions, developing adaptive strategies for selecting $\delta$, and extending CPGR to dynamic and streaming graph settings.
+Appendix A Appendix
+[…]
+Figure 9: Neighborhood tree of vertex $u_{2}\in U$ that shows the path $\omega$ taken from root to the vertex $w_{j}\in W$ at the leaf and number of edges $d_{u_{2},\omega}$ for each node using the tuple $(\omega,d_{u_{2},\omega})$.
+A.1 FM algorithm
 In Algorithm 3 we give a high-level description of the FM algorithm.
-The FM algorithm first constructs neighborhood trees for each vertex $u_{i}$ in $U$, which are labeled binary trees of depth $r$, whose nodes are labeled by a bit string $\omega$, where $0\leq|\omega|\leq r$.
+The FM algorithm first constructs neighborhood trees (Figure 9) for each vertex $u_{i}$ in $U$, which are labeled binary trees of depth $r$, whose nodes are labeled by a bit string $\omega$, where $0\leq|\omega|\leq r$.
 The root node of a neighborhood tree contains all the vertices in $W$ and is labeled by the empty string, $\epsilon$.
@@
 Each node in the neighborhood tree of $u_{i}$ stores $d_{i,\omega}$, the number of neighbors of $u_{i}$ in the set $W_{\omega}$ of vertices of $W$ that are associated with the node with label $\omega$, i.e., $d_{i,\omega}=|N(u_{i})\cap W_{\omega}|$, where $N(u_{i})$ is the set of the neighbors of $u_{i}$.
-In Figure 1(b), we show the neighborhood tree of vertex $u_{2}$ for the bipartite graph with $|U|=|W|=n=8$ given in Figure 1(a).
+In Figure 9, we show the neighborhood tree of vertex $u_{2}$ for the bipartite graph with $|U|=|W|=n=8$ given in Figure 2.
 Vertex $w_{3}$ is a leaf and would have $\omega=010$ in all neighborhood trees of the graph.
@@
 5:  $\hat{k}\leftarrow\bigg\lfloor{\frac{\delta\log n}{\log(2n^{2}/\hat{m})}}\bigg\rfloor$
-6:     Starting with $t=1$ and $U_{t}=U$, iterate while $t\leq\hat{k}$ to select $\hat{k}$ vertices based on $c_{0}$ and $c_{1}$ (calculated as in Equation A.1) to form the ordered set $K_{q}=\{y_{1},y_{2},\ldots,y_{t}\}\subseteq W$.
+6:   Starting with $t=1$ and $U_{t}=U$, iterate while $t\leq\hat{k}$ to select $\hat{k}$ vertices based on $c_{0}$ and $c_{1}$ (calculated as in Equation 3) to form the ordered set $K_{q}=\{y_{1},y_{2},\ldots,y_{t}\}\subseteq W$.
 At each stage $t$, obtain set $U_{t+1}=\left\{u\in U_{t}\mid y_{1},\ldots,y_{t}\in N(u)\right\}$ which omits vertices from $U$ that are not adjacent to $K_{q}$.
@@
 10:For each clique in $\mathcal{C}=\{C_{1},\ldots,C_{q}\}$ add edges from all the vertices in the left partition $U_{q}$ to an additional vertex $z_{q}$, and from $z_{q}$ to all the vertices in the right partition $K_{q}$.
-This would compress the graph by replacing $|U_{q}|\times|K_{q}|$ edges with $|U_{q}|+|K_{q}|$ edges.
-11:The remaining edges in $G$ are trivial cliques and are added in the compressed graph $G^{*}$.
-12:Output: $G^{*}$, the compressed graph of $G$.
+This would restructure the graph by replacing $|U_{q}|\times|K_{q}|$ edges with $|U_{q}|+|K_{q}|$ edges.
+11:The remaining edges in $G$ are trivial cliques and are added in the restructured graph $G^{*}$.
+12:Output: $G^{*}$, the restructured graph of $G$.
 The algorithm computes $\hat{k}$ and then, in Line 5, it performs $\hat{k}$ iterations (indexed by $t$) to select vertices for the right partition of clique $C_{q}$.
-In each iteration it selects vertices at leaf nodes of the neighborhood trees by following a path based on $d_{i,\omega}$ of child nodes and the number of distinct ordered subsets of the neighbourhood tree at each level.
+In each iteration it selects vertices at leaf nodes of the neighborhood trees by following a path based on $d_{i,\omega}$ of child nodes and the number of distinct ordered subsets of the neighborhood tree at each level.
 In order to select a path at each level, the algorithm calculates $c_{0}$ and $c_{1}$ as follows:
-(A.1) $c_{j}=C\left(U_{t},\mathscr{C}_{t,\omega\cdot j}\right)=\sum_{i:{u_{i}\in U_{t}}}d_{i,\omega\cdot j}\cdot\left(d_{i}-1\right)^{[{k}-t]},\,j=0,1.$
+$c_{j}=C\left(U_{t},\mathscr{C}_{t,\omega\cdot j}\right)=\sum_{i:{u_{i}\in U_{t}}}d_{i,\omega\cdot j}\cdot\left(d_{i}-1\right)^{[{k}-t]},\,j=0,1.$ (3)
 $c_{j}$ counts the number of ordered sets in $\mathscr{C}_{t,\omega}$ whose elements are adjacent to $u_{i}$, where $\mathscr{C}_{t,\omega}$ denotes the collection of all ordered sets $K$ of $W$ of size $k$ in iteration $t$, with $t\leq k$.
@@
 The edges that are not removed in this process form trivial cliques.
-The FM algorithm compresses the graph by adding a new vertex set $Z$ to the bipartite graph, thus converting it into a tripartite graph (Figure 1(c)), where each vertex in $Z$ corresponds to one clique extracted by the algorithm.
+The FM algorithm restructures the graph by adding a new vertex set $Z$ to the bipartite graph, thus converting it into a tripartite graph (Figure 2a), where each vertex in $Z$ corresponds to one clique extracted by the algorithm.
 Each vertex of the left and right partitions of $\delta$-clique $C_{q}$ is then connected via an edge to a new vertex $z_{q}\in Z$, thus forming a tripartite graph.
-This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus compressing the graph.
-The compressed graph obtained by the algorithm preserves the path information of the original graph.
+This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus reducing the number of edges in the graph.
+The restructured graph obtained by the algorithm preserves the path information of the original graph.
 The running time of the FM algorithm is $O(mn^{\delta}\log^{2}n)$.
-The FM algorithm can be extended to the case of non-bipartite graphs, where it computes a compression of a graph in time $O(mn^{\delta}\log^{2}n)$, as shown by Feder and Motwani [9].
-A.2 Example: CPGC Execution.
-We now show how CPGC works on a bipartite graph $G$ with partitions $U$ and $W$, $n=|U|=|W|=8$, and $|E|=54$, shown in Figure 6(a).
-CPGC (Algorithm 1), first initializes $q=0,n=|W|=|U|=8,\hat{m}=|E|=54,\mathcal{C}=\emptyset$, and $d_{w}=[0]_{n}$.
-It then calculates the degree of each vertex $d_{w_{j}}$ in Lines 2-4, as shown in Table 1.
-It calculates $\hat{k}(n,m,\delta)=2$ in Line 5.
-Thus, the condition for the while loop in Line 6 is met and CPGC calls CSA in Line 7.
+The FM algorithm can be extended to the case of non-bipartite graphs, where it restructures a graph in time $O(mn^{\delta}\log^{2}n)$, as shown by Feder and Motwani [14].
+A.2 Example: CPGR Execution
+We now show how CPGR works on a bipartite graph $G$ with partitions $U$ and $W$, $n=|U|=|W|=8$, and $|E|=54$, shown in Figure 10(a).
+CPGR (Algorithm 1), first initializes $q=0,n=|W|=|U|=8,\hat{m}=|E|=54,\mathcal{C}=\emptyset$, and $d_{w}=[0]_{n}$.
+It then calculates the degree of each vertex $d_{w_{j}}$ in Lines 6-8, as shown in Table 3.
+It calculates $\hat{k}(n,m,\delta)=2$ in Line 9.
+Thus, the condition for the while loop in Line 10 is met and CPGR calls CSA in Line 11.
 With the given inputs, CSA sorts $d_{w}$ in non-increasing order.
 Let $\{w_{4},w_{2},w_{3},w_{5},w_{6},w_{1},w_{7},w_{8}\}$ be the non-increasing order with corresponding degrees given as $d_{w_{\pi(j)}}$ in Table 3.
-CPGC step $j$ 1 2 3 4 5 6 7 8
+CPGR step $j$ 1 2 3 4 5 6 7 8
 Initialization $d_{w_{j}}$ 6 7 7 8 7 7 6 6
@@
 After $1^{st}$iteration $d_{w_{j}}$ 6 0 0 1 0 7 6 6
-Table 1: Degrees of vertices in partition $W$ at different steps of CPGC.
+Table 3: Degrees of vertices in partition $W$ at different steps of CPGR.
 $u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$
@@
 $u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$u_{6}$$u_{7}$$u_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$$w_{8}$$z_{1}$$z_{2}$
-Figure 6: Clique partitioning: (a) given bipartite graph $G(U,W)$; (b) extracted clique $C_{1}$; (c) extracted clique $C_{2}$; (d) graph with trivial cliques; and (e) compressed tripartite graph $G(U,Z,W)$
+Figure 10: Clique partitioning: (a) given bipartite graph $G(U,W)$; (b) extracted clique $C_{1}$; (c) extracted clique $C_{2}$; (d) graph with trivial cliques; and (e) restructured tripartite graph $G^{*}(U,Z,W)$
 In Line 5, CSA forms set $\mathcal{K}$, where $\mathcal{K}=\{w_{4},w_{2},w_{3},w_{5},w_{6}\}$ has all vertices whose degrees are greater than or equal to $d_{w_{\pi(\hat{k})}}=7$.
 This results in $\gamma=2$ in Line 7, therefore CSA extracts two cliques in Lines 8-13.
-Since $q$ is initialized to 0 in CPGC, the for loop in Line 8 iterates two times for $c=1$ and $2$.
+Since $q$ is initialized to 0 in CPGR, the for loop in Line 8 iterates two times for $c=1$ and $2$.
 In Line 9, CSA forms $K_{1}=\{w_{4},w_{2}\}$, and in Line 10, it forms the left partition $U_{K_{1}}$ of bipartite clique $\mathcal{C}_{1}$, by selecting the common neighbors of the vertices in $K_{1}$.
@@
 It then forms the set of bipartite cliques extracted in the current execution in Line 14, $\mathscr{C}=\{C_{1},C_{2}\}$.
-The two cliques are shown in Figures 6(b) and 6(c).
-CSA returns $(\mathscr{C},\hat{A},\hat{d}_{w})$ and then CPGC updates the set $\mathscr{C}=\{C_{1},C_{2}\}$, $d_{w}$, the adjacency matrix $A$, and the number of edges $\hat{m}=54-28=26$, in Lines 8-10.
-In Line 12, CPGC updates $q=2$ and finally, in Line 13 it updates $\hat{k}$ according to the new value of $\hat{m}$ in Line 11 which results in $\hat{k}=1$.
-Since $\hat{k}=1$, it means CPGC extracts only trivial bipartite cliques (shown in Figure 6(d)) which do not contribute to graph compression.
-Thus, it does not meet the condition in the while loop (Line 6) and CPGC terminates.
-CPGC then compresses the graph by adding two vertices, $z_{1}$, and $z_{2}$, corresponding to the two cliques $C_{1}$ and $C_{2}$, and adds the corresponding edges to form the tripartite graph.
+The two cliques are shown in Figures 10(b) and 10(c).
+CSA returns $(\mathscr{C},\hat{A},\hat{d}_{w})$ and then CPGR updates the set $\mathscr{C}=\{C_{1},C_{2}\}$, $d_{w}$, the adjacency matrix $A$, and the number of edges $\hat{m}=54-28=26$, in Lines 12-15.
+In Line 16, CPGR updates $q=2$ and finally, in Line 17 it updates $\hat{k}$ according to the new value of $\hat{m}$ in Line 15 which results in $\hat{k}=1$.
+Since $\hat{k}=1$, it means CPGR extracts only trivial bipartite cliques (shown in Figure 10(d)) which do not contribute to reduction in the number of edges in the restructured graph.
+Thus, it does not meet the condition in the while loop (Line 6) and CPGR terminates.
+CPGR then restructures the graph by adding two vertices, $z_{1}$, and $z_{2}$, corresponding to the two cliques $C_{1}$ and $C_{2}$, and adds the corresponding edges to form the tripartite graph.
 The edges in the given graph $G(U,W,E)$, that are not part of any $\delta$-clique are connected directly as shown in Figure 10(e).
-A.3 Proofs
-Proof of Theorem 4.1.
-  • Proof.
-    Let’s assume that CPGC extracts only one $\delta$-clique $C_{q}$ from the given graph $G$.
-In the $\delta$-clique $C_{q}$, the right partition $K_{q}\subseteq W$ is formed by selecting $\hat{k}$ vertices from $W$ and the left partition $U_{q}\subseteq U$ is formed with the common neighbors of $K_{q}$, forming a complete bipartite graph with $E_{q}$ edges.
-The compressed graph $G^{*}$ is formed with the same left and right partitions $U$ and $W$ of the graph $G$, and a third partition $Z$ which is a set of additional vertices associated with each of the $\delta$-cliques that CPGC extracts.
-Our main concern is with the edges $(u_{i},w_{j})\in E^{*}$ in the compressed graph $G^{*}$.
-$E^{*}$ contains the edges that replace the $\delta$-clique $C_{q}$ by adding the additional vertex $z_{q}\in Z$, and the set of edges $\hat{E}=E\setminus E_{q}$ which were not part of the $\delta$-clique $C_{q}$.
-It can be easily seen that $\hat{E}\subset E$, that is, the edges in $\hat{E}$ are edges in $G$, and the remaining edges $E^{*}\setminus\hat{E}$ in $E^{*}$ are the edges that replace $C_{q}$ in $G^{*}$.
-Each edge $(u_{i},w_{j})\in E_{q}$, is replaced by two edges, $(u_{i},z_{q})$ and $(z_{q},w_{j})$, where $u_{i}\in U$, $w_{j}\in W$, and $z_{q}\in Z$.
-Therefore, for each edge $(u_{i},w_{j})\in E_{q}$, there exists a path from $u_{i}$ to $w_{j}$ composed of two edges, $(u_{i},z_{q})$ and $(z_{q},w_{j})$, that passes through the additional vertex $z_{q}\in Z$, thus preserving the path information.
-This holds true for all the $\delta$-cliques extracted by CPGC.
-Proof of Lemma 4.1.
-  • Proof.
-    In the $\delta$-clique based graph compression performed by CPGC, the compressed graph $G^{*}(U,W,Z,E^{*})$ is obtained from $G$ by replacing $|E_{q}|=|U_{q}\times K_{q}|$ edges in a $\delta$-clique $C_{q}$ with $|E^{*}|=|U_{q}|+|K_{q}|$ edges in $G^{*}$ and adding an additional vertex $z_{q}\in Z$ for each extracted $\delta$-clique $C_{q}$.
-The size of the right partition of $C_{q}$ is $|K_{q}|={k}=k(n,m,\delta)=\Big{\lfloor}{\frac{\delta\log n}{\log(2n^{2}/m)}}\Big{\rfloor}$.
-If $m<2n^{2-\frac{\delta}{2}}$ then the size of the right partition $|K_{q}|$ is $k=1$.
-Therefore, the number of edges in the $\delta$-clique is $|E_{q}|=|U_{q}|\cdot 1=|U_{q}|$.
-Those edges are replaced by $|E^{*}|=|U_{q}|+\hat{k}=|U_{q}|+1$ edges in $G^{*}$.
-Thus, replacing the $\delta$-clique in $G$ with $m<2n^{2-\frac{\delta}{2}}$, the number of edges in $G^{*}$ would actually increase by 1, which does not lead to a compression of $G$.
-A.4 Extension to non-bipartite graphs.
-CPGC can be extended to compress non-bipartite graphs using a similar technique to that described in [9].
-Consider a directed graph $H(V,E)$ with $n$ vertices and $m$ edges.
-Initially, we transform $H$ into a bipartite graph $G\left(L,R,E^{\prime}\right)$, where each vertex $v\in V$ is duplicated into $v_{L}\in L$ and $v_{R}\in R$.
-For each directed edge $(u,v)$ in $E$, we create a directed edge $\left(u_{L},v_{R}\right)$ in $E^{\prime}$, with the direction from $L$ to $R$.
-Furthermore, for each vertex $v\in V$ we add a directed edge from $v_{R}$ to $v_{L}$.
-The proposed compression algorithm is then applied to the bipartite graph $G$, which includes only the edges from $L$ to $R$.
-Once the compressed graph is computed we add to it the $n$ directed edges from $u_{R}$ to $u_{L}$, corresponding to each vertex $u\in V$.
-Since the path information of the original graph is preserved by this transformation, it allows different graph algorithms to work on the compressed graph as well.
-Undirected general graphs can be transformed first into directed graphs by simply replacing each undirected edge $(u,v)$ in the original graph by two directed edges $(u,v)$ and $(v,u)$, and then applying the technique presented above to transform the directed graph into a bipartite graph.
-The time complexity and the compression ratio of CPGC are the same as in the case of the bipartite graphs.
-A similar approach for converting a general graph to a bipartite graph and then compressing the obtained bipartite graph was employed in [3, 5].